### v1

# Improving Complex Reasoning over Knowledge Graph with Logic-Aware Curriculum Tuning

Abstract
Answering complex logical queries over incomplete knowledge graphs (KGs) is challenging. Most previous works have focused on learning entity/relation embeddings and simulating first-order logic operators with various neural networks. However, they are bottlenecked by the inability to share world knowledge to improve logical reasoning, thus resulting in suboptimal performance. In this paper, we propose a complex logical reasoning schema over knowledge graphs upon large language models (LLMs), containing a curriculum-based logical-aware instruction tuning framework, named LACT. Specifically, we augment the arbitrary first-order logical queries via binary tree decomposition, to stimulate the reasoning capability of LLMs. To address the difficulty gap among different types of complex queries, we design a simple and flexible logic-aware curriculum learning framework. Experiments across widely used datasets demonstrate that LACT has substantial improvements (brings an average +5.5% MRR score) over advanced methods, achieving the new state-of-the-art. Our code and model will be released at GitHub and huggingface soon.
1 Introduction
Large-scale knowledge graphs (KGs) such as FreeBase  (?) stores structural knowledge in a collection of fact triplets and are widely adopted by many domains. Unfortunately, KGs are often incomplete, leaving many missing triplets undiscovered. Thus, complex logical reasoning over such KGs  (??) is challenging and has attracted much attention in the recent years. A complex logical query can be represented with First-Order Logic (FOL) that includes logical operators such as conjunction ()($\wedge$)( ∧ ), disjunction ()($\vee$)( ∨ ), negation (¬)($\neg$)( ¬ ), and existential quantifier ()($\exists$)( ∃ ), etc. A more direct approach involves the representation of computation graphs as Directed Acyclic Graphs (DAGs), which can be resolved through the systematic traversal of Knowledge Graphs (KG). This process entails the allocation of suitable entities to intermediate variables based on their structural attributes (?).
Figure 1: Schematic illustration: a) Answering logical query over KG with LACT. b) The framework of Logic-Aware Curriculum Tuning over LLama. We leverage binary tree decomposition strategy (Seen in Section 4.2) to construct logic-rich FT corpus and Curriculum learning strategy (Seen in Section 4.3) to fine-tune a base LLM. c) Performing reasoning using well-designed prompts.
Inspired by the success of knowledge graph embedding (KGE) (??), a line of research proposes to answer complex logical queries by learning query embedding and simulating logical operators with well-designed neural networks  (????) Current research based on embeddings primarily focuses on the creation of diverse latent space geometries, such as vectors (?), boxes (?), hyperboloids (?), and probabilistic distributions (?), to effectively capture the semantic position and logical coverage of knowledge graph entities.
However, these approaches are limited in their performance due to the following. (1) Limited information: The information contained in a knowledge graph is usually incomplete and limited. When only the information from the knowledge graph can be used, it is difficult to answer some complex reasoning that lacks relevant information. (2) High complexity of logical queries: The intricacies of world knowledge determine the complexity of reasoning in practical applications, which determines that it is difficult to model the relationship of world knowledge through simple space geometries figures that may lose potentially complex relationship information (?), thus limiting the effect of complex logical reasoning. (3) Generalizability: KGE method for a particular KG can not generalize to other KGs which limits the applicability of these approaches in real-world scenarios where KGs can vary widely in terms of their structure and content.
Recently, large language models (LLMs)  (???) show outperforming capabilities to a wide range of tasks (?????).  (??) construct prompt templates and apply LLMs as text-generators to answer complex queries. However, LLM without fine-tuning suffers from hallucination problem (?), especially when faced with such a knowledge-intensive task that generates answers on an incomplete KG rather than simple retrieval. Besides, previous tasks relied on manual classification of queries to improve performance, which is unrealistic in large-scale practical applications and also limits the types of queries. Finally, Previous methods generally decompose the problem into sub-problems, which greatly increases the cost of reasoning, especially considering that previous methods are generally based on closed-source models such as GPT. Therefore, compared to pure prompt engineering, we prefer to fine-tune our own model to solve the above problems.
In this paper, we propose Logic-Aware Curriculum Tuning (LACT  [Uncaptioned image]), a novel fine-tune framework for answering complex logical query, which stimulates the ability of LLMs to perform complex reasoning on knowledge graphs. We propose a strategy to incorporate the knowledge contained in the KGs into our training corpus to activate the corresponding knowledge of the LLMs and supplement the missing relevant knowledge of the LLMs during the fine-tuning process. At the same time, we have proven that data argument by binary tree decomposition can arouse the corresponding capabilities of LLMs and effectively improve their reasoning performance. At last, we show that curriculum learning(?) can effectively smooth the difficulty differences between different types of queries and greatly improve the results of difficult queries. In summary, our contribution is three-folded:
  • We propose a logic-aware curriculum fine-tuning (LACT) paradigm for complex logical reasoning over KGs.
  • LACT achieves state-of-the-art performance beyond embedding-based and PLM-based methods, using only a 7B llama2 model.
  • Through extensive experimentation, we found that fine-tuning corpus constructed with rigorous logical context over KGs and curriculum learning can significantly enhance LLM logical reasoning ability.
2 Related Works
2.1 Logical Reasoning over Knowledge Graph
Knowledge graphs can be widely used to enhance natural language reasoning (??), understanding (??), and generation (??). Given a FOL query over a KG, complex logical reasoning aims to answer correct entities, which contains not only multi-hop paths but logical operators(??). Most of current approaches concentrated on learning meaningful query embeddings (?????). Neuralizing logical operators through a specific embedding space, thereby embedding FOL queries into a vector space(???), or probabilistic distribution(?), and predict answers by locating nearest neighbours to answer set representation. Additionally, approaches such as CQD(?) have focused on improving the performance of complex reasoning tasks through the answer composition of simple intermediate queries, and QTO (?) proposes query computation tree optimization that can efficiently find the exact optimal solutions. However, embedding-based methods usually lack interpretability as there is no explicit mapping between the embedding and the set of entities, and this limits their generalization ability to more complicated query structures.
Moreover, PLM-based methods consider complex logical reasoning as text-based tasks by modeling pre-trained language models. The short textual descriptions are organized as an input sequence and encoded by the PLMs (???). However, limited by the performance limitations of the base model, the PLM method has always been deficient in understanding world knowledge and reasoning capabilities, limiting its performance in complex reasoning.
2.2 LLMs for KG Reasoning
In recent years, substantial advancements have been witnessed in the domain of LLMs (????). Among these, instruction tuning (IT) (?) and the alignment of the model (?) with human preferences stand out.
Within the realm of LLM, the integration of LLMs with Knowledge Graphs (KG) (???) constitutes a prominent and consequential research avenue.
Leveraging its potent generative capabilities, LLMs prove invaluable in addressing Knowledge Graph-related tasks, including but not limited to Knowledge Graph Completion (KGC) (??), entity alignment (?), Knowledge Graph Question Answering (KGQA) (?), and others (?). Consequently, the synergy between Knowledge Graphs for LLMs (KG4LLM) and LLMs for Knowledge Graphs (LLM4KG) emerges as an essential focal point, bearing significance in advancing the collective capabilities of both entities.
We focus on applying LLMs in the Complex Logical Reasoning task, which has not been carefully studied yet. (?) made the initial attempt by prompt engineer but it lacks in-depth research and simply uses LLMs as text generators.
2.3 Curriculum Learning
The idea of training a neural network in an easy-to-difficult fashion can be traced back to  (??).Based on these works, Curriculum Learning is first proposed in (?). They designed several toy experiments to demonstrate the benefits of curriculum strategy both in image classification and language modeling. When focusing on the field of NLP, by experimenting with several heuristics,  (??) migrated the success of CL to NLU tasks. (??) improved the machine translation modeling by carefully designing different curriculum. Recently, with the rise of LLM, (?) discovered the huge potential of CL in in-context learning, while (?) focus on the improvement of CL for LLM in pre-train. However, we are committed to exploring the huge potential of CL in fine-tuning LLM.
3 Preliminary
3.1 Knowledge Graph
In our work, a knowledge graph is $\mathcal{G}=(\mathcal{E,R,T})$ where $\mathcal{E,R}$ are the set of entity and relation respectively. With regard to generality, KG can be organized as a set of triplets $\{\mathcal{T}=(v_{s},r,v_{t})|v_{s},v_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $v_{s}/v_{t}$ denotes the head/tail entity.
3.2 Complex logical query
Complex logical query is used for retrieving and manipulating data stored in knowledge graphs, which is grounded in a subset of FOL. The process of answering a complex logical query involves trying to match a suitable results using the composition of queries:
$q[v_{?}]=\exists v:q_{1}\wedge q_{2}\wedge\cdots\wedge q_{n},$ (1)
or,
$q[v_{?}]=\exists v:q^{\prime}_{1}\vee q^{\prime}_{2}\vee\cdots\vee q^{\prime}_{n},$ (2)
where $q$ denotes a FOL query. Note that Eq. (1) is conjunctive normal form (CNF) and Eq. (2) is disjunctive normal form (DNF). The two can be equivalently converted to each other via De Morgan’s law. Following previous works (?), we focus on modeling the operations: projection $r(\cdot)$, conjunction ($\wedge$), disjunction ($\vee$), and negation ($\neg$). Additionally, note that existential positive first-order (EPFO) queries only includes projection, conjunction ($\wedge$), and disjunction ($\vee$).
4 Methodology
4.1 Instruction Tuning on LLMs
In this section, we introduce how to incorporate the KG information in the text-based prompt.
When applying LLMs to complex logical reasoning, we denote an LLM as $\mathcal{M}$ which is a text decoder to generate the corresponding output. If we start from the above definition, this task can be modelled as a text generation task. However, triplet generation is different from vanilla text generation because the entities and the relation in the triplet prompt have complex semantic information defined by the given KG. In fact, we want the generated answers to be entities that exist in KG itself. Without these knowledge, the predicted answers are unreliable and unstable. Thus, incorporating the KG information into the prompt to provide more auxiliary information is the key to engage LLMs in complex logical reasoning.
In particular, when we fine-tune $\mathcal{M}$, we can treat the training corpus as a set of question-answer pairs $(\mathcal{S},\mathcal{A})$. For the task of complex logical reasoning over knowledge graph, the input textual sequence $\mathcal{S}$ consists of the description of question $\mathcal{D}$, knowledge graph neighbourhood information(i.e. related triplets) $\mathcal{X}$ and logical query. In our work, we used a simple but effective method called greedy depth traversal algorithm to search for neighbourhood information and we simply discarded all samples that exceeded the token limit. (Detailed algorithm and token distribution can be found in Appendix A). The logical query contains the textual information about the query $q_{\tau}$ with the query structure $\tau$ and specific query content $Q_{\tau}$ that needs to be processed, which can be denoted as $f_{1}(q_{\tau})$. Likewise,we can denote the output answer A as $f_{2}(V_{\tau})$, where $f_{2}$ indicates textualization of $V_{\tau}$ here. In summary, the fine-tune training corpus $\mathcal{C}$ can be expressed in the following form:
$\mathcal{C}=(\mathcal{S},\mathcal{A})=(\mathcal{D}\oplus\mathcal{X}\oplus f_{1}(q_{\tau}),f_{2}(V_{\tau})).$ (3)
The model $\mathcal{M}$(parameterized by $\theta$) is fine-tuned by the next token prediction task. We fine-tune $\mathcal{M}$ to obtain the final model by maximizing the log-likelihood of the next token. The training objective can be formulated as
$\mathcal{L}=-\frac{1}{\left|\mathcal{C}\right|}\sum_{i=1}^{\left|\mathcal{C}\right|}\log P_{\mathcal{M}}\left(c_{i}\mid c_{<i}\right),$ (4)
where $c_{i}(i=1,2,...,\left|\mathcal{C}\right|)$ represents the textual tokens of the training corpus $\mathcal{C}$. For our task, the training objective can be transferred as
$\mathcal{L}=-\frac{1}{\left|\mathcal{C}\right|}\sum_{i=1}^{\left|\mathcal{C}\right|}\log P_{\mathcal{M}}\left(\mathcal{A}\mid\mathcal{S}\right).$ (5)
4.2 Data Augmentation by Binary Tree Decomposition
This section introduces how to build fine-tuning corpora that make LLMs (Large Language Models) logic-aware based on instruction tuning.
Chain of thought (COT) enables models to decompose multi-step problems into intermediate steps, subsequently improving the reasoning abilities of LLMs (?). However, pure-prompt based reasoning needs more incontext memory to perform complex logical reasoning. Considering complex logical queries, which query structure can be transferred into the form of a DAG and its hierarchical structure becomes a natural fit for decomposition into a series of sub-problems. So we propose a method for data augment based on Binary Tree Decomposition Mechanism to stimulate LLMs with the potential to decompose a complex query into a chain of simple queries. Binary Tree Decomposition Mechanism. The Binary Tree Decomposition Mechanism is divided into the following three steps:
Query Computation Tree. For a complex FOL query, like the example shown in Figure 1, its computation graph that is a directed acyclic graph can be converted into a tree where the root node is $v_{?}$. query computation tree. The answer variable and the constant entities in the query correspond to root and leaf nodes in the query computation tree. Each edge in the query computation tree points from the child node to the parent node. It can be recursively deduced that the subtree rooted at any non-leaf node in the tree corresponds to a subquery. In Appendix B, we provide a systematical procedure for transforming an FOL query to its query computation tree.
Binary Tree Decomposition. For the one-to-many intersection/union structures in the tree, we separate each parent node into two child nodes. Note that the union branches merging step may create one-to-many structures that consist of both intersection and union edges, take Figure 1 for an example. This can be taken care of by first separating $v_{?}$ into an intersection structure ($v_{3}^{\prime}$ and $v_{5}^{\prime}$ in the example), and then separating the child node into an intersection structure ($v_{1}^{\prime}$ and $v_{2}^{\prime}$ for $v_{3}^{\prime}$, $v_{4}^{\prime},v_{3}$ for $v_{5}^{\prime}$), where $v^{\prime}$ denotes an intermediate entity retrieved by Neighborhood Retrieval Algorithm (Seen in Appendix A).
Reverse Level Traversal. Finally, we decompose the binary computation tree into independent branches. Since the root node of the calculation tree is the answer entity, we perform a hierarchical traversal of all non-leaf nodes of the binary tree in reverse. As shown in Figure 1, the complex FOL query is decomposed into a sequence:
$[(v_{1},r,v_{1}^{\prime})$ $,(v_{2},r,v_{2}^{\prime}),(v_{3},r,v_{5}^{\prime}),(v_{4},r,v_{4}^{\prime}),$
$(v_{1}^{\prime},r,v_{3}^{\prime})$ $,\wedge,(v_{2}^{\prime},r,v_{3}^{\prime}),\neg,(v_{4}^{\prime},r,v_{5}^{\prime}),\wedge,(v_{3},r,v_{5}^{\prime}),$
$(v_{3}^{\prime},r,v_{?})$ $,\wedge,(v_{5}^{\prime},r,v_{?})].$
Data Augmentation. Now we can turn any loopless FOL query into a series of separate subqueries. We use a defined template to integrate the decomposition process into the answers to the training corpus. So, the training corpus $\mathcal{C}$ can be transferred into the following form:
$\mathcal{C}=(\mathcal{S},\mathcal{A})=(\mathcal{D}\oplus\mathcal{X}\oplus f_{1}(q_{\tau}),f_{2}(V_{\tau,\mathcal{D}ecomposed})),$ (6)
where $V_{\tau,\mathcal{D}ecomposed}$ indicates the answer corresponding to the logical query with the decomposition reasoning path.
4.3 Fine-tuning Enhanced by Curriculum Learning
As mentioned in previous sections, though decomposing into chain responses, complex queries still vary greatly in difficulty and complexity due to differences between query structure.
Naturally, we believe that these different types of samples should not be simply lumped together. Intuitively, we incorporate curriculum learning into our training. To be specific, In view of the particularity of complex reasoning data, when we decompose it into logical chains, naturally, we can use the number of decomposed sub-logical queries as a natural difficulty discriminator to select different types of queries, e.g., a 1p query would be defined as difficulty-1, while a 2p query, which can be decomposed into two projection queries and an intersection query, would be defined as difficulty-3. The detailed difficulty discriminating process will be shown in Appendix (Table S1).
Finally, we divided samples into three parts: easy samples, medium samples and difficult samples according to the difficulty level. Correspondingly, our training process is also divided into three stages. After we did some exploratory experiments, we did not simply train three data sets in the order of easy-medium-difficult. On the contrary, we decided to first use 80% easy samples, 10% medium samples, and 10% difficult samples for the first stage of training and the subsequent two-stage training process is a Leto, and experimental results in the next few sections also proved that this is effective.
Method $avg_{p}$ $avg_{ood}$ $avg_{n}$ 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni FB15K GQE 28.0 20.1 - 54.6 15.3 10.8 39.7 51.4 27.6 19.1 22.1 11.6 - - - - - Query2Box 38.0 29.3 - 68.0 21.0 14.2 55.1 66.5 39.4 26.1 35.1 16.7 - - - - - BetaE 41.6 34.3 11.8 65.1 25.7 24.7 55.8 66.5 43.9 28.1 40.1 25.2 14.3 14.7 11.5 6.5 12.4 CQD-CO 46.9 35.3 - 89.2 25.3 13.4 74.4 78.3 44.1 33.2 41.8 21.9 - - - - - CQD-Beam 58.2 49.8 - 89.2 54.3 28.6 74.4 78.3 58.2 67.7 42.4 30.9 - - - - - ConE 49.8 43.4 14.8 73.3 33.8 29.2 64.4 73.7 50.9 35.7 55.7 31.4 17.9 18.7 12.5 9.8 15.1 GNN-QE 72.8 68.9 38.6 88.5 69.3 58.7 79.7 83.5 69.9 70.4 74.1 61.0 44.7 41.7 42.0 30.1 34.3 QTO 74.0 71.8 49.2 89.5 67.4 58.8 80.3 83.6 75.2 74.0 76.7 61.3 61.1 61.2 47.6 48.9 27.5 LARK 56.1 43.1 18.4 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 82.6 71.9 56.9 93.5 73.5 59.6 92.3 82.3 76.8 75.9 74.6 60.4 81.2 61.6 52.0 43.5 41.7 FB15K-237 GQE 16.3 10.3 - 35.0 7.2 5.3 23.3 34.6 16.5 10.7 8.2 5.7 - - - - - Query2Box 20.1 15.7 - 40.6 9.4 6.8 29.5 42.3 21.2 12.6 11.3 7.6 - - - - - BetaE 20.9 14.3 5.5 39.0 10.9 10.0 28.8 42.5 22.4 12.6 12.4 9.7 5.1 7.9 7.4 3.5 3.4 CQD-CO 21.8 15.6 - 46.7 9.5 6.3 31.2 40.6 23.6 16.0 14.5 8.2 - - - - - CQD-Beam 22.3 15.7 - 46.7 11.6 8.0 31.2 40.6 21.2 18.7 14.6 8.4 - - - - - FuzzQE 24.0 17.4 7.8 42.8 12.9 10.3 33.3 46.9 26.9 17.8 14.6 10.3 8.5 11.6 7.8 5.2 5.8 ConE 23.4 16.2 5.9 41.8 12.8 11.0 32.6 47.3 25.5 14.0 14.5 10.8 5.4 8.6 7.8 4.0 3.6 GNN-QE 26.8 19.9 10.2 42.8 14.7 11.8 38.3 54.1 31.1 18.9 16.2 13.4 10.0 16.8 9.3 7.2 7.8 QTO 33.5 27.6 15.5 49.0 21.4 21.2 43.1 56.8 38.1 28.0 22.7 21.4 16.8 26.7 15.1 13.6 5.4 LARK 50.7 41.0 10.6 73.6 40.5 26.8 46.1 43.1 49.9 22.9 62.8 28.3 6.5 3.4 23.2 16.5 3.2 LACT 57.0 44.4 21.9 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 NELL995 GQE 18.6 12.5 - 32.8 11.9 9.6 27.5 35.2 18.4 14.4 8.5 8.8 - - - - - Query2Box 22.9 15.2 - 42.2 14.0 11.2 33.3 44.5 22.4 16.8 11.3 10.3 - - - - - BetaE 24.6 14.8 5.9 53.0 13.0 11.4 37.6 47.5 24.1 14.3 12.2 8.5 5.1 7.8 10.0 3.1 3.5 CQD-CO 28.8 20.7 - 60.4 17.8 12.7 39.3 46.6 30.1 22.0 17.3 13.2 - - - - - CQD-Beam 28.6 19.8 - 60.4 20.6 11.6 39.3 46.6 25.4 23.9 17.5 12.2 - - - - - FuzzQE 27.0 18.4 7.8 47.4 17.2 14.6 39.5 49.2 26.2 20.6 15.3 12.6 7.8 9.8 11.1 4.9 5.5 ConE 27.2 17.6 6.4 53.1 16.1 13.9 40.0 50.8 26.3 17.5 15.3 11.3 5.7 8.1 10.8 3.5 3.9 GNN-QE 28.9 19.6 9.7 53.3 18.9 14.9 42.4 52.5 30.8 18.9 15.9 12.6 9.9 14.6 11.4 6.3 6.3 QTO 32.9 24.0 12.9 60.7 24.1 21.6 42.5 50.6 31.3 26.5 20.4 17.9 13.8 17.9 16.9 9.9 5.9 LARK 52.9 26.9 12.4 87.8 45.7 33.5 51.3 48.7 23.1 22.2 20.6 41.1 9.9 5.9 24.5 13.3 7.3 LACT 60.1 32.0 17.2 91.4 53.6 40.6 62.2 54.9 31.4 34.8 27.0 34.0 16.0 21.2 21.0 16.3 11.6
Table 1: Test MRR results (%) on complex query answering across all query types. $avg_{p}$ is the average on EPFO queries; $avg_{ood}$ is the average on out-of-distribution (OOD) queries, which consist pi/ip/2u/ip queries; $avg_{n}$ is the average on queries with negation.
4.4 Reasoning Module
We use the final LACT LLM as the answer generator, as shown in Figure 1(c). We retrieve relevant information and textualize the FOL query, and finally we populate it into the template in 4.4 to generate responses.
We use the LLM to do a simple text generation task to get the answer. After fine-tune, LACT LLM can follow the output mode in training stage in Figure 1, so we can extract final answers through simple regular expressions with the template in 4.4.
Prompt 1: Query Prompt Template of LACT. [Uncaptioned image]
Prompt 2: Answer Template of LACT. [Uncaptioned image]
5 Experiments
5.1 Training Settings
Training datasets We opt for the the most popular datasets: FB15K, FB15K-237, NELL995. Detailed information about dataset is listed in  D.1. We used the training set of the above dataset as the original training data.
Training Details We use open-source model LLaMA-2-base, including two different parameter sizes: 7B and 13B, as the base model for fine-tuning. All LLaMA-2-7B and LLaMA-2-13B models are trained by fully fine-tuning.
For the fully fine-tuning setting, we use the AdamW optimizer to train the model with 1 epoch and the batch size is 128. We use 8 NVIDIA A100 GPUS for training with the learning rate of 3e-6.
Training Cost We trained for a total of approximately 10 hours using 8*A100 during three training stages.
5.2 Experimental Settings
Baseline Methods For comparing with KGE, we chose the following representative methods as baselines: GQE (?), Query2Box(Q2B) (?), BetaE (?), CQD (?), ConE (?), GNN-QE (?), QTO (?). We also compared our method to an LLM-based method, LARK  (?).
Evaluation Protocol Following previous works (?), we use mean reciprocal rank (MRR) as standard evaluation protocols for evaluating complex reasoning over knowledge graph. In the filtered setting, all easy and hard answers are filtered out during ranking. The detail could be found in Appendix  D.3.
5.3 Main Results
The main experiment results of three datasets are shown in Table 1. The baselines were trained on 1p/2p/3p/2i/3i queries, hence other than these 4 types of EPFO queries serve as OOD queries, and we report the average result on these queries in $avg_{ood}$. We observe that LACT significantly outperforms baseline methods across all datasets. Notably, LACT yields an averag gain of 7.3%, 2.9%, and 6.3% on $avg_{p}$, $avg_{odd}$, and $avg_{n}$, compared to the previous SOTA method, especially more challenging datasets like FB15K-237 and NELL995. This suggests that our method has better reasoning capability and captures a broad range of relations to effectively utilize this capability for enhancing the performance of complex queries.
Figure 2: Results of Ablation Studies.(a) Comparison PPL and train loss results of whether to use BTD based on FB15k.(b) Comparison PPL results of whether use CL based on FB15K high-difficulty queries.
Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2 67.2 42.3 38.3 61.6 44.8 34.1 36.9 44.2 28.4 44.7 38.5 36.9 32.1 30.0    + IT 94.6 68.8 60.2 84.5 66.7 56.0 60.2 69.5 42.3 66.5 54.4 41.0 38.8 36.9     +BTD 91.5 72.3 65.6 89.7 75.2 60.1 65.4 72.5 49.9 74.3 66.4 48.9 46.5 42.5
Table 2: Accuracy results (%) of whether to use BTD on hard complex query answering across all query types, evaluated on FB15k.
BTD CL 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2 w/IT 94.6 68.8 60.2 84.5 66.7 56.0 60.2 69.5 42.3 66.5 54.4 41.0 38.8 36.9 94.2 72.3 65.6 89.7 75.2 60.1 65.4 72.5 49.9 74.3 66.4 48.9 46.5 42.5 94.6 70.9 61.3 84.4 72.7 58.2 63.2 69.5 47.3 68.5 64.0 42.3 40.9 37.1 94.8 78.7 69.2 94.8 88.1 79.3 80.5 80.7 67.1 90.6 70.4 59.3 53.6 46.7
Table 3: Accuracy results (%) of whether to use CL on hard complex query answering across all query types, evaluated on FB15k.
Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2-7B 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 Mistral-7B 82.6 59.5 34.2 59.6 58.5 59.1 39.1 59.2 33.7 20.4 39.2 31.4 22.4 13.5 Qwen1.5-7B 81.3 57.2 36.8 62.9 58.1 63.9 38.6 60.4 31.8 20.7 41.9 28.7 20.8 11.6
Table 4: MRR results (%) based on differnet base models across all query types, evaluated on FB15k-237.
Figure 3: The results of the main experiment. We evaluate the performance of three current state-of-the-art methods on three datasets.
5.4 Ablation Studies
To verify the effectiveness of the LACT design, we conduct a two-part ablation study. The first part is designed to verify the effectiveness of logical chain decomposition and the second part is designed to verify the effectiveness of curriculum learning.
Effect of Binary Tree Decomposition (BTD). As shown in Figure 2, logical chain decomposition can stimulate LLM’s ability of logical decomposition, thereby greatly improving the performance of difficult queries.
From a training perspective, as shown in Figure 2, although perplexity (PPL) and training loss of decomposed queries before training was slightly higher than that of ordinary queries, we found that as training progresses, the loss and PPL of decomposed queries will quickly decrease to levels much lower than ordinary queries, proving that chain decomposition is effective to reduce the difficulty of learning complex queries.
Effect of Curriculum Learning. Curriculum learning, as illustrated in Table 3, greatly alleviates the gap between difficult training samples and the understanding ability of LLMs.
We can observe from Figure 2 that compared with random shuffle sequence training, difficult training samples under curriculum learning gradually become easier to understand. It is worth mentioning that we found that the gain of curriculum learning on training corpus that has not been decomposed by logical chains is very small, which supports our theory from the side. It is difficult for LLMs to understand the difficulty difference between undecomposed samples, so curriculum learning is also difficult to take effect.
Figure 4: Performance of scaling LACT on FB15K-237 with different a) model and b) data scales.
5.5 Transferability Study
Considering the diversity of complex reasoning tasks, we can divide transferability into two levels, task-level and dataset-level transferability.
Task-level transferability. The results in Table 1 show that our method achieves a relative gain of 9.9% on the OOD task, which demonstrates the strong generalization of our fine-tuning framework. Even in the OOD queries, as shown in Table 5, more than 95% of test samples can still follow logical chain reasoning. These phenomena indicate strong generalization ability of LACT.
Metric pi ip 2u up Prodecomposed 98.7 100.0 97.8 100.0 Protrue,decomposed 98.6 99.9 97.8 99.6
Table 5: In OOD queries, the proportion of queries that can be decomposed and the proportion of queries that can be decomposed correctly on fb15k.
Figure 5: Ablation experimental results of Accuracy (%) trained on FB15k and tested on FB15K-237, compared to models trained on all mixed training data.
Dataset-level transferability. In fact, almost all KGE methods, even if some of the optimization methods claim not to require training, require a KGE base model adapted to a specific dataset, which leads to the inherent defect of extremely poor Transferability of the KGE method. However, as previous research has shown, fine-tuning of LLMs is mainly to stimulate the knowledge and capabilities of potentially relevant downstream tasks contained in LLM pre-training. This has also become the theoretical basis for the transferability of fine-tuning methods for LLMs. The results in Figure 5 show that the reasoning ability stimulated by one dataset can still be demonstrated in another dataset, which reflects well in the query performance which only dropped less than 5%.
Model-level transferability. We tried analytical experiments with different base models to determine whether our LACT was universal. Obviously, all types of queries have been improved to a certain extent due to the progress of the base model. Experimental results show that our LACT is suitable for different pedestal models and has strong generalization.
5.6 Scalability Study
For verifying the scalability of LACT, we scale LACT to different model sizes and data volumes.
Performance on different model size. We tried scaling model size to see if LACT would have an impact when operating on a larger scale. As Figure 4 shows, the performance of our method improves as the model size increases.
Performance on different data size. We conducted experiments on different ratios of training data to verify the robustness of LACT.
Figure 6: (left) MRR performance of LACT and previous SOTA methods at different difficulties based on FB15K. (right) The correlation between related information completeness and accuracy evaluated on FB15K-237, we selected 3p query and pi query with the same inference path length as the task types. We assume that the completeness of all simple queries is 1.
6 Discussion
Figure 7: Inference results of ChatGPT and LACT on 2i query case, respectively.
6.1 When and Where Does LACT Work?
The performance of LACT would be related to the following two aspects: I. The completeness of relevant information extracted from KG. II. Sophistication of complex reasoning.
LACT performs consistently better with more complete information. We take the form of a posteriori that set the completeness of relevant triplets to the proportion of triplets in the inference path of complex reasoning in the provided context, (For example, if the inference path of a 2i query is (Turing Award, winner, Yoshua Bengio), (Canada, citizen,Yoshua Bengio), and we can only retrieve (Turing Award, winner, Yoshua Bengio) in incomplete KG, then our completeness of relevant triplets to 1/2.) and set the completeness of simple queries that can be directly inferred to 1, to obtain the relation between Accuracy and correlation information completeness. As seen in Figure  6, LACT obtains a significant gain when the completeness of relevant information increased, though, with zero relevant information, it remains a certain amount of complex reasoning ability.
LACT performs consistently better on higher difficulties As mentioned before, we simply divide the difficulty of the query by the number of hops in the query. The results in Figure  6 show that our model yields more gain in tasks of higher-level difficulty and complexity, which benefits from our unique and sophisticated fine-tuning framework.
6.2 Case Study
To have a close look, we perform the case studies by analyzing the results of LACT and ChatGPT (GPT-3.5-turbo-0613). As shown in Figure 7, ChatGPT cannot make good use of incomplete knowledge graphs for reasoning in some cases. Conversely, LACT performs reasoning through a complete logical chain, making maximum use of the relevant information provided and deducing the correct answer, which greatly improves the reasoning ability.
7 Conclusion
In this paper, we present a simple and effective fine-tuning framework LACT to boost the complex logical reasoning ability over KGs. LACT is a two-stage method that consists of both Binary Tree Decomposition and Curriculum Learning and can be applied to various size LLMs with different data sizes. We empirically demonstrate the effectiveness and universality of the LACT on a series of widely-used knowledge graph datasets. Further analyses reveal the underlying mechanism of our method, and investigate When and Why Does LACT Work. We hope that our work can inspire more research on combining knowledge graphs and LLMs.
Appendix
Appendix A Neighborhood Retrieval Algorithm
A.1 Retrieval Algorithm
To strike a balance between the completeness of relevant information and the token number limit of LLMs, we search for as many relevant triplets as possible along the possible paths.
Particularly, for the 1p query, we simply find all the triplets containing the entity or the relation.
For another query, as shown in Figure S1 for each leaf node in DAG, we do depth traversal on the graph. For each step in the traversal process, if this step is a projection, we search for all the possible triplets. Otherwise, we perform corresponding operations on intersection and union respectively to filter out the corresponding entities.
We continue this traversal until the obtained entity is empty or reaches the root node. All triplets during the traversal are related to triplets.
A.2 Over-limit Solutions and Token distribution
In all experimental setups, we used the Max Seq Length of 4096, and we simply discarded all out-of-bounds samples and recorded them as 0 at the time of evaluation. In fact, after using our information retrieval algorithm, most of our samples were controlled below 4096. As shown in Figure S2, compared with the algorithm of searching all possible related triples, our retrieval algorithm greatly reduces the consumption of tokens.
Figure S1: A case on Neighborhood Retrieval Algorithm
Figure S2: Probability distribution of the number of tokens in FB15K and all-context Retrieval Algorithm.
1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Number of subqueries 1 2 3 3 5 4 4 3 4 3 5 4 4 4 Difficulty 1 1 2 2 3 3 3 2 3 2 3 3 3 3
Table S1: Difficulty of different query types where 1 means easy, 2 means medium, 3 means hard.
Figure S3: Query structures, illustrated in their query computation graph representations
Appendix B Conversion Between FOL Expression and Query Computation Tree
The conversion from an FOL expression (disjunctive normal form) to its query computation tree involves two steps: dependency graph generation, and union branches duplication.
Dependency Graph Generation. Upon encountering a First-Order Logic (FOL) expression, our primary procedure entails the allocation of distinct nodes to individual variables, while assigning a unique node to the constant entity within each one-hop atom. It is important to acknowledge that multiple nodes may represent the same constant entity, given its occurrence in various one-hop atoms. Subsequently, undirected edges are employed to establish connections between nodes in accordance with the defined one-hop atoms. Specifically, if $e^{i}_{j}=r(v^{\prime},v)$ (or $r(c,v)$), then we connect the nodes of $v^{\prime}$ (or $c$) and $v$ by an edge $r_{i}$. Similarly, if $e^{i}_{j}=\lnot r(v^{\prime},v)$ (or $\lnot r(c,v)$), then we connect the nodes of $v^{\prime}$ (or $c$) and $v$ by an edge $\lnot r_{i}$. The variable $i$ serves as a distinguishing label for edges emanating from distinct conjunctions. The formulated undirected dependency multigraph must conform to a tree structure, signifying a connected and acyclic graph. Choosing the node $v_{?}$ as the root, we establish edge directions ensuring that they uniformly point from child nodes to their respective parent nodes, with due consideration to handling inverse relations. Notably, constant entities inherently function as leaf nodes within the tree, given that each entity node exclusively connects to a single variable node.
Union Branches Duplication Then we handle the duplication branches in the query computation tree. On the path $\tau$ from root to every leaf node, if exists, we find the first node $v_{i}$ such that the edges between $v_{i}$ and its child node $v_{j}$ are all of the same relations, but in different conjunctions: $r_{t_{1}},r_{t_{2}},\dots,r_{t_{p}}$. We merge these edges into a single edge $r_{t_{1},t_{2},\dots t_{p}}$, since they all correspond to the same one-hop atom but in different conjunctions, they can be merged by the distributive law:
$(A\land B)\lor(A\land C)\Leftrightarrow A\land(B\lor C)$ (7)
We assert that there is a subpath from $v_{i}$ to some $v_{k}$ within the path $\tau$ that only consists of edges $r_{t_{1},t_{2},\dots t_{p}}$, and $v_{k}$ is connected to different child nodes by relations from conjunctions $t_{1},t_{2},\dots t_{p}$. These edges are annotated as unions, while the remaining one-to-many structures are designated as intersections. Consequently, the multigraph transforms into a simple graph, devoid of multiple edges.
Appendix C Difficulty
We divide the difficulty by the number of decomposed subqueries. Query types and their corresponding difficulties are shown in Table S1.
Appendix D Experiment Details
D.1 Dataset Details
$\bullet$ FB15K is based on Freebase, a large collaborative knowledge graph project that was created by Google. FB15k contains about 15,000 entities, 1,345 relations, and 592,213 triplets (statements that assert a fact about an entity).
$\bullet$ FB15K-237 is a subset of FB15k, containing 14,541 entities, 237 relations, and 310,116 triplets. The relations in FB15k-237 are a subset of the relations in FB15k, and were created to address some of the limitations of FB15k, such as the presence of many irrelevant or ambiguous relations, and to provide a more challenging benchmark for knowledge graph completion models.
$\bullet$ NELL995 was created using the Never-Ending Language Learning (NELL) system, which is a machine learning system that automatically extracts knowledge from the web by reading text and inferring new facts. NELL995 contains 9,959 entities, 200 relations, and 114,934 triplets. The relations in NELL995 cover a wide range of domains, including geography, sports, and politics.
D.2 Query Structure
For a fair comparison, we use the 14 types of complex queries generated by the same rules in (?). The query structure of each type is shown in Figure S3.
For each complex query, answers are divided into easy answers and hard answers, based on whether answer can be derived by existing edges in the graph directly. Specifically, in the valid/test set, the easy answers are the entities that can be inferred by edges in the training/valid graph, while hard answers are those that need to be inferred by predicting missing edges in the valid/test graph. Referring to previous work, we calculate standard evaluation metrics including mean reciprocal rank (MRR), in the filtered setting where all easy and hard answers are filtered out during ranking.
D.3 Evaluation Protocol Detail
For each complex query, its answers are divided into easy answers and hard answers, based on whether the answer can be derived by existing edges in the graph directly. Specifically, in the valid/test set, the easy answers are the entities that can be inferred by edges in the training/valid graph, while hard answers are those that need to be inferred by predicting missing edges in the valid/test graph. Referring to previous work (?), we calculate standard evaluation metrics including mean reciprocal rank (MRR), in the filtered setting where all easy and hard answers are filtered out during ranking.

### v4

@@
 Abstract
-Answering complex logical queries over incomplete knowledge graphs (KGs) is challenging.
+Answering complex queries over incomplete knowledge graphs (KGs) is a challenging task.
 Most previous works have focused on learning entity/relation embeddings and simulating first-order logic operators with various neural networks.
 However, they are bottlenecked by the inability to share world knowledge to improve logical reasoning, thus resulting in suboptimal performance.
-In this paper, we propose a complex logical reasoning schema over knowledge graphs upon large language models (LLMs), containing a curriculum-based logical-aware instruction tuning framework, named LACT.
+In this paper, we propose a complex reasoning schema over KG upon large language models (LLMs), containing a curriculum-based logical-aware instruction tuning framework, named LACT.
 Specifically, we augment the arbitrary first-order logical queries via binary tree decomposition, to stimulate the reasoning capability of LLMs.
@@
 Experiments across widely used datasets demonstrate that LACT has substantial improvements (brings an average +5.5% MRR score) over advanced methods, achieving the new state-of-the-art.
-Our code and model will be released at GitHub and huggingface soon.
-1 Introduction
-Large-scale knowledge graphs (KGs) such as FreeBase  (?) stores structural knowledge in a collection of fact triplets and are widely adopted by many domains.
+Codehttps://github.com/TianleXia0914/LACT
+Introduction
+Industrial-scale knowledge graphs (KGs) like FreeBase store structural knowledge in a collection of fact triplets and are widely adopted by many domains.
 Unfortunately, KGs are often incomplete, leaving many missing triplets undiscovered.
-Thus, complex logical reasoning over such KGs  (??) is challenging and has attracted much attention in the recent years.
-A complex logical query can be represented with First-Order Logic (FOL) that includes logical operators such as conjunction ()($\wedge$)( ∧ ), disjunction ()($\vee$)( ∨ ), negation (¬)($\neg$)( ¬ ), and existential quantifier ()($\exists$)( ∃ ), etc.
-A more direct approach involves the representation of computation graphs as Directed Acyclic Graphs (DAGs), which can be resolved through the systematic traversal of Knowledge Graphs (KG).
-This process entails the allocation of suitable entities to intermediate variables based on their structural attributes (?).
+Thus, complex logical reasoning over such incomplete KGs (Hamilton et al.
+2018; Bai et al.
+2023) is a challenging task and has attracted much attention in the recent years.
+In the realm of complex logical queries, Existential First Order queries with a single free variable ($\bm{EFO_{1}}$) emerge as a powerful tool that includes logical operators such as the basic projection operation ()($\exists$)( ∃ ) and the corresponding negation operation (¬)($\neg$)( ¬ ), two relational operation intersection ()($\wedge$)( ∧ ) and union ()($\vee$)( ∨ ), etc.
+A very straightforward approach is to represent the computation graph as a Directed Acyclic Graph (DAG), which can be addressed by a systematic traversal of a knowledge graph (KG).
+This process entails the allocation of suitable entities to intermediate variables based on their structural attributes (Dalvi and Suciu 2007).
 Figure 1: Schematic illustration: a) Answering logical query over KG with LACT.
 b) The framework of Logic-Aware Curriculum Tuning over LLama.
-We leverage binary tree decomposition strategy (Seen in Section 4.2) to construct logic-rich FT corpus and Curriculum learning strategy (Seen in Section 4.3) to fine-tune a base LLM.
+We leverage the binary tree decomposition strategy (Seen in Section Methodology) to construct a logic-rich FT corpus and the Curriculum learning strategy (Seen in Section Methodology) to fine-tune a base LLM.
 c) Performing reasoning using well-designed prompts.
-Inspired by the success of knowledge graph embedding (KGE) (??), a line of research proposes to answer complex logical queries by learning query embedding and simulating logical operators with well-designed neural networks  (????) Current research based on embeddings primarily focuses on the creation of diverse latent space geometries, such as vectors (?), boxes (?), hyperboloids (?), and probabilistic distributions (?), to effectively capture the semantic position and logical coverage of knowledge graph entities.
-However, these approaches are limited in their performance due to the following.
-(1) Limited information: The information contained in a knowledge graph is usually incomplete and limited.
-When only the information from the knowledge graph can be used, it is difficult to answer some complex reasoning that lacks relevant information.
-(2) High complexity of logical queries: The intricacies of world knowledge determine the complexity of reasoning in practical applications, which determines that it is difficult to model the relationship of world knowledge through simple space geometries figures that may lose potentially complex relationship information (?), thus limiting the effect of complex logical reasoning.
-(3) Generalizability: KGE method for a particular KG can not generalize to other KGs which limits the applicability of these approaches in real-world scenarios where KGs can vary widely in terms of their structure and content.
-Recently, large language models (LLMs)  (???) show outperforming capabilities to a wide range of tasks (?????).
-(??) construct prompt templates and apply LLMs as text-generators to answer complex queries.
-However, LLM without fine-tuning suffers from hallucination problem (?), especially when faced with such a knowledge-intensive task that generates answers on an incomplete KG rather than simple retrieval.
+In the wake of knowledge graph embedding (KGE)’s resounding success, (Bordes et al.
+2013; Bai et al.
+2021), a series of work seeks to answer complex logical queries by learning query embedding and simulating logical operators with well-designed neural networks  (Chen, Hu, and Sun 2022; Zhu et al.
+2022; Zhang et al.
+2021; Arakelyan et al.
+2020).
+In contemporary research, the exploration of KGE method predominantly centers around the intricate design of various embedding representations, including geometric embedding like Euclidean geometry modeling (Hamilton et al.
+2018; Ren, Hu, and Leskovec 2019) and Riemannian geometry modeling (Choudhary et al.
+2021), and embedding based on probability modeling (Ren, Hu, and Leskovec 2019), seeking to encapsulate the semantic representation and vector mapping of KGs’ entities and relations.
+However, the embedding-based approach described above also has some limitations.
+(1) Limited information: Knowledge graphs often contain incomplete and insufficient information.
+When only the information from the knowledge graph is available, it becomes challenging to perform complex reasoning that requires additional or missing information.
+(2) Insufficient utilization of semantic information: Current knowledge graph embedding (KGE) methods primarily focus on the structural information of the knowledge graph, paying little attention to the actual semantics of entities and relationships.
+(3) High complexity of logical queries: he intricate nature of real-world knowledge contributes to the complexity of reasoning in practical applications.
+As a result, it is challenging to accurately model the relationships within real-world knowledge using simple geometric representations.
+Such simplifications may overlook potentially complex relational information (Choudhary et al.
+2021), thereby limiting the effectiveness of advanced logical reasoning.
+(4) Generalizability: A Knowledge Graph Embedding (KGE) model tailored to a specific knowledge graph cannot generalize across different knowledge graphs.
+This limitation hinders the practical application of such approaches in real-world scenarios, where knowledge graphs often vary significantly in both structure and content.
+Recently, large language models (LLMs)  (Achiam et al.
+2023; Touvron et al.
+2023a) have showed outperforming capabilities for a wide range of tasks (Ouyang et al.
+2022; Zhong et al.
+2023b; Peng et al.
+2023; Lu et al.
+2023; Guo et al.
+2025).
+With the rise of this trend, (Choudhary and Reddy 2023; Liu et al.
+2024a) construct prompt templates and apply LLMs as text-generators to answer complex queries.
+However, LLM without fine-tuning suffers from hallucination problem (Zhang et al.
+2023c), especially when faced with such a knowledge-intensive task that generates answers on an incomplete KG rather than simple retrieval.
 Besides, previous tasks relied on manual classification of queries to improve performance, which is unrealistic in large-scale practical applications and also limits the types of queries.
 Finally, Previous methods generally decompose the problem into sub-problems, which greatly increases the cost of reasoning, especially considering that previous methods are generally based on closed-source models such as GPT.
-Therefore, compared to pure prompt engineering, we prefer to fine-tune our own model to solve the above problems.
+Therefore, compared to pure prompt engineering, we prefer to fine-tune our model to solve the above problems.
 In this paper, we propose Logic-Aware Curriculum Tuning (LACT[Uncaptioned image]), a novel fine-tune framework for answering complex logical query, which stimulates the ability of LLMs to perform complex reasoning on knowledge graphs.
@@
 At the same time, we have proven that data argument by binary tree decomposition can arouse the corresponding capabilities of LLMs and effectively improve their reasoning performance.
-At last, we show that curriculum learning(?) can effectively smooth the difficulty differences between different types of queries and greatly improve the results of difficult queries.
-In summary, our contribution is three-folded:
-  • We propose a logic-aware curriculum fine-tuning (LACT) paradigm for complex logical reasoning over KGs.
-  • LACT achieves state-of-the-art performance beyond embedding-based and PLM-based methods, using only a 7B llama2 model.
-  • Through extensive experimentation, we found that fine-tuning corpus constructed with rigorous logical context over KGs and curriculum learning can significantly enhance LLM logical reasoning ability.
-2 Related Works
-2.1 Logical Reasoning over Knowledge Graph
-Knowledge graphs can be widely used to enhance natural language reasoning (??), understanding (??), and generation (??).
-Given a FOL query over a KG, complex logical reasoning aims to answer correct entities, which contains not only multi-hop paths but logical operators(??).
-Most of current approaches concentrated on learning meaningful query embeddings (?????).
-Neuralizing logical operators through a specific embedding space, thereby embedding FOL queries into a vector space(???), or probabilistic distribution(?), and predict answers by locating nearest neighbours to answer set representation.
-Additionally, approaches such as CQD(?) have focused on improving the performance of complex reasoning tasks through the answer composition of simple intermediate queries, and QTO (?) proposes query computation tree optimization that can efficiently find the exact optimal solutions.
-However, embedding-based methods usually lack interpretability as there is no explicit mapping between the embedding and the set of entities, and this limits their generalization ability to more complicated query structures.
-Moreover, PLM-based methods consider complex logical reasoning as text-based tasks by modeling pre-trained language models.
-The short textual descriptions are organized as an input sequence and encoded by the PLMs (???).
+At last, we show that curriculum learning (Bengio et al.
+2009) can effectively smooth the difficulty differences between different types of queries and greatly improve the results of difficult queries.
+In summary, our contribution manifests in three distinct facets:
+  • We propose a Logic-Aware Curriculum tuning (LACT) paradigm for complex logical reasoning over KGs.
+  • LACT achieves state-of-the-art performance beyond embedding-based and PLM-based methods, using only 7B models.
+  • Through extensive experimentation, we found that SFT Corpus constructed by queries on incomplete KGs via Binary Tree Decomposition and Curriculum Learning can significantly enhance LLM logical reasoning ability.
+Related Works
+Logical Reasoning over Knowledge Graph
+Knowledge graphs can be widely used to enhance natural language reasoning (Zhang and Yao 2022; Zhao et al.
+2023), understanding (Zhong et al.
+2023a; Liu et al.
+2023), and generation (Ding, Wang, and Liu 2023; Pan et al.
+2024b).
+Given a $EFO_{1}$ query over a KG, complex logical reasoning aims to answer the correct entities, which contains both multi-hop queries and logical operators(Guu, Miller, and Liang 2015; Hamilton et al.
+2018) over a incomplete KG.
+Most of the current approaches have focused on learning meaningful KG embeddings including (Arakelyan et al.
+2020; Chen, Hu, and Sun 2022; Zhang et al.
+2021; Zhu et al.
+2022; Wang et al.
+2023b).
+Neuralizing logical operators through a specific embedding space, thereby embedding $EFO_{1}$ queries into a multi-dimensional vector space(Hamilton et al.
+2018; Ren, Hu, and Leskovec 2019), or into a specific probability distribution(Ren, Hu, and Leskovec 2019; Choudhary et al.
+2021), then obtain the final answer set by fitting the nearest neighbour representation based on relational operations.
+Additionally, recent work like CQD (Arakelyan et al.
+2020) its optimization work (Arakelyan et al.
+2024) improves performance by working to reduce query difficulty by decomposing complex queries into simple one-hop queries, QTO (Bai et al.
+2023) introduces query computation tree optimization and LMPNN (Wang et al.
+2023f) aggregates logical messages passed on the query graph by training an MLP network.
+Some work (Yin, Wang, and Song 2023) has begun to expand into the area of cyclic query.
+Despite their efficacy, embedding-based methods often suffer from a lack of generalization due to their specificity in embedding knowledge graphs.
+Meanwhile, this limitation on the embedding range hinders their ability to effectively generalize to more complex query structures.
+Moreover, PLM-based methods view complex logical reasoning as text-generation tasks by modeling pre-trained language models.
+$EFO_{1}$ queries are organized into input-output sequence pairs after textualization and encoded by PLM (Wang et al.
+2023c; Xu et al.
+2023; Wang et al.
+2023c).
 However, limited by the performance limitations of the base model, the PLM method has always been deficient in understanding world knowledge and reasoning capabilities, limiting its performance in complex reasoning.
-2.2 LLMs for KG Reasoning
-In recent years, substantial advancements have been witnessed in the domain of LLMs (????).
-Among these, instruction tuning (IT) (?) and the alignment of the model (?) with human preferences stand out.
-Within the realm of LLM, the integration of LLMs with Knowledge Graphs (KG) (???) constitutes a prominent and consequential research avenue.
-Leveraging its potent generative capabilities, LLMs prove invaluable in addressing Knowledge Graph-related tasks, including but not limited to Knowledge Graph Completion (KGC) (??), entity alignment (?), Knowledge Graph Question Answering (KGQA) (?), and others (?).
+LLMs for KG Reasoning
+In recent years, substantial advancements have been witnessed in the domain of LLMs (Achiam et al.
+2023; Touvron et al.
+2023b; Peng et al.
+2023; Zhong et al.
+2023b; Zeng et al.
+2022).
+Among these, instruction tuning (IT) (Ouyang et al.
+2022) and the alignment of the model (Wang et al.
+2023e) with human preferences stand out.
+Within the realm of LLM, the integration of LLMs with Knowledge Graphs (KG) (Pan et al. 2024a; Wang et al. 2023a; Luo et al. 2024) constitutes a prominent and consequential research avenue.
+Leveraging its potent generative capabilities, LLMs prove invaluable in addressing Knowledge Graph-related tasks, including but not limited to Knowledge Graph Completion (KGC) (Zhu et al.
+2023; Zhang et al.
+2023b), entity alignment (Zhang et al.
+2023a), Knowledge Graph Question Answering (KGQA) (Luo et al.
+2024), and others (Luo et al.
+2023).
 Consequently, the synergy between Knowledge Graphs for LLMs (KG4LLM) and LLMs for Knowledge Graphs (LLM4KG) emerges as an essential focal point, bearing significance in advancing the collective capabilities of both entities.
-We focus on applying LLMs in the Complex Logical Reasoning task, which has not been carefully studied yet.
-(?) made the initial attempt by prompt engineer but it lacks in-depth research and simply uses LLMs as text generators.
-2.3 Curriculum Learning
-The idea of training a neural network in an easy-to-difficult fashion can be traced back to  (??).Based on these works, Curriculum Learning is first proposed in (?).
-They designed several toy experiments to demonstrate the benefits of curriculum strategy both in image classification and language modeling.
-When focusing on the field of NLP, by experimenting with several heuristics,  (??) migrated the success of CL to NLU tasks.
-(??) improved the machine translation modeling by carefully designing different curriculum.
-Recently, with the rise of LLM, (?) discovered the huge potential of CL in in-context learning, while (?) focus on the improvement of CL for LLM in pre-train.
+Our research endeavours center around the utilization of Language Models (LLMs) for the Complex Logical Reasoning task, an area that remains relatively unexplored.
+(Choudhary and Reddy 2023) made the initial attempt by prompt engineer but it lacks in-depth research and simply uses LLMs as text generators.
+Curriculum Learning
+The concept of progressively training neural networks from simple to complex configurations has its origins dating back to (Elman 1993; Krueger and Dayan 2009).
+Based on these works, Curriculum Learning is first proposed in (Bengio et al.
+2009).
+A series of illustrative experiments were meticulously crafted to showcase the advantages of employing curriculum strategies in both image classification and language modeling.
+When focusing on the field of NLP, by experimenting with several heuristics,  (Sachan and Xing 2016; Xu et al.
+2020) migrated the success of CL to NLU tasks.
+(Ding et al.
+2021; Zhou et al.
+2021) improved the machine translation modeling by carefully designing different curricula.
+Recently, with the rise of LLM, (Liu et al.
+2024b) discovered the huge potential of CL in in-context learning, while (Wang et al.
+2023d) focus on the improvement of CL for LLM in pre-train.
 However, we are committed to exploring the huge potential of CL in fine-tuning LLM.
-3 Preliminary
-3.1 Knowledge Graph
+Preliminary
+Knowledge Graph
 In our work, a knowledge graph is $\mathcal{G}=(\mathcal{E,R,T})$ where $\mathcal{E,R}$ are the set of entity and relation respectively.
-With regard to generality, KG can be organized as a set of triplets $\{\mathcal{T}=(v_{s},r,v_{t})|v_{s},v_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $v_{s}/v_{t}$ denotes the head/tail entity.
-3.2 Complex logical query
-Complex logical query is used for retrieving and manipulating data stored in knowledge graphs, which is grounded in a subset of FOL.
-The process of answering a complex logical query involves trying to match a suitable results using the composition of queries:
+With regard to generality, KG can be considered as a collection of triplets $\{\mathcal{T}=(v_{s},r,v_{t})|v_{s},v_{t}\in\mathcal{E},r\in\mathcal{R}\}$, where $v_{s}/v_{t}$ denotes the head/tail entity.
+Complex logical query
+Complex logical query is used for retrieving and manipulating data stored in knowledge graphs, which is grounded in a subset of $EFO_{1}$ query.
+The process of answering a complex logical query involves trying to match suitable results using the composition of queries:
 $q[v_{?}]=\exists v:q_{1}\wedge q_{2}\wedge\cdots\wedge q_{n},$ (1)
@@
 $q[v_{?}]=\exists v:q^{\prime}_{1}\vee q^{\prime}_{2}\vee\cdots\vee q^{\prime}_{n},$ (2)
-where $q$ denotes a FOL query.
+where $q$ denotes a $EFO_{1}$ query and $q_{i}$ denotes a sub-query.
 Note that Eq.
-(1) is conjunctive normal form (CNF) and Eq.
-(2) is disjunctive normal form (DNF).
+(1) is a conjunctive normal form (CNF) and Eq.
+(2) is a disjunctive normal form (DNF) (Ren, Hu, and Leskovec 2019).
 The two can be equivalently converted to each other via De Morgan’s law.
-Following previous works (?), we focus on modeling the operations: projection $r(\cdot)$, conjunction ($\wedge$), disjunction ($\vee$), and negation ($\neg$).
-Additionally, note that existential positive first-order (EPFO) queries only includes projection, conjunction ($\wedge$), and disjunction ($\vee$).
-4 Methodology
-4.1 Instruction Tuning on LLMs
-In this section, we introduce how to incorporate the KG information in the text-based prompt.
-When applying LLMs to complex logical reasoning, we denote an LLM as $\mathcal{M}$ which is a text decoder to generate the corresponding output.
-If we start from the above definition, this task can be modelled as a text generation task.
-However, triplet generation is different from vanilla text generation because the entities and the relation in the triplet prompt have complex semantic information defined by the given KG.
+Methodology
+Instruction Tuning on LLMs
+Within this section, we elucidate the methodology for seamlessly integrating Knowledge Graph (KG) information into the textual prompt.
+In the context of complex logical reasoning, we designate a Language Model (LLM) as $\mathcal{M}$ serving as a text decoder to produce the corresponding output.
+By commencing with the aforementioned definition, we can frame this task as a text generation problem.
+n contrast to vanilla text generation, triplet generation involves more complexity due to the intricate semantic information associated with the entities and relations in the triplet prompt, as defined by the given knowledge graph (KG).
 In fact, we want the generated answers to be entities that exist in KG itself.
-Without these knowledge, the predicted answers are unreliable and unstable.
-Thus, incorporating the KG information into the prompt to provide more auxiliary information is the key to engage LLMs in complex logical reasoning.
+Without this knowledge, the predicted answers are unreliable and unstable.
+Thus, to engage LLM in complex logical reasoning, a crucial step involves incorporating knowledge graph (KG) information into the prompt.
+This integration provides additional auxiliary context, effectively enhancing the performance of LLM.
 In particular, when we fine-tune $\mathcal{M}$, we can treat the training corpus as a set of question-answer pairs $(\mathcal{S},\mathcal{A})$.
-For the task of complex logical reasoning over knowledge graph, the input textual sequence $\mathcal{S}$ consists of the description of question $\mathcal{D}$, knowledge graph neighbourhood information(i.e.
+When considering complex reasoning over KGs, the input textual sequence $\mathcal{S}$ consists of the description of question $\mathcal{D}$, knowledge graph neighbourhood information(i.e.
 related triplets) $\mathcal{X}$ and logical query.
 In our work, we used a simple but effective method called greedy depth traversal algorithm to search for neighbourhood information and we simply discarded all samples that exceeded the token limit.
-(Detailed algorithm and token distribution can be found in Appendix A).
+(Detailed algorithm and token distribution can be found in Appendix).
 The logical query contains the textual information about the query $q_{\tau}$ with the query structure $\tau$ and specific query content $Q_{\tau}$ that needs to be processed, which can be denoted as $f_{1}(q_{\tau})$.
@@
 We fine-tune $\mathcal{M}$ to obtain the final model by maximizing the log-likelihood of the next token.
-The training objective can be formulated as
+The training loss can be formulated as
 $\mathcal{L}=-\frac{1}{\left|\mathcal{C}\right|}\sum_{i=1}^{\left|\mathcal{C}\right|}\log P_{\mathcal{M}}\left(c_{i}\mid c_{<i}\right),$ (4)
@@
 $\mathcal{L}=-\frac{1}{\left|\mathcal{C}\right|}\sum_{i=1}^{\left|\mathcal{C}\right|}\log P_{\mathcal{M}}\left(\mathcal{A}\mid\mathcal{S}\right).$ (5)
-4.2 Data Augmentation by Binary Tree Decomposition
+Data Augmentation by Binary Tree Decomposition
 This section introduces how to build fine-tuning corpora that make LLMs (Large Language Models) logic-aware based on instruction tuning.
-Chain of thought (COT) enables models to decompose multi-step problems into intermediate steps, subsequently improving the reasoning abilities of LLMs (?).
-However, pure-prompt based reasoning needs more incontext memory to perform complex logical reasoning.
+Chain of thought (COT) enables models to decompose multi-step problems into intermediate steps, subsequently improving the reasoning abilities of LLMs (Wei et al.
+2022).
+However, pure-prompt-based reasoning needs more in-context memory to perform complex logical reasoning.
 Considering complex logical queries, which query structure can be transferred into the form of a DAG and its hierarchical structure becomes a natural fit for decomposition into a series of sub-problems.
@@
 Query Computation Tree.
-For a complex FOL query, like the example shown in Figure 1, its computation graph that is a directed acyclic graph can be converted into a tree where the root node is $v_{?}$.
-query computation tree.
-The answer variable and the constant entities in the query correspond to root and leaf nodes in the query computation tree.
-Each edge in the query computation tree points from the child node to the parent node.
-It can be recursively deduced that the subtree rooted at any non-leaf node in the tree corresponds to a subquery.
-In Appendix B, we provide a systematical procedure for transforming an FOL query to its query computation tree.
+Existing embedding-based methods rely on DNF to avoid embedding divergence (Ren and Leskovec 2020).
+Our method bypasses logical operator learning.
+Consequently, this allows us to convert any query into a logic-equivalent tree form through iterative sub-query recursion.
+For a complex $EFO_{1}$ query, like the example shown in Figure 1, its computation graph that is a directed acyclic graph can be converted into a tree where the root node is $v_{?}$.
+The tail entities and intermediate entities in complex queries are in one-to-one correspondence with the root nodes and leaf nodes in the generated calculation tree respectively.
+In view of the reverse order that the tail entity brings to the root node, the edges in the tree corresponding to the relationship in each query are child-to-parent directions.
+In the Appendix, We demonstrate a simple but systematic method for converting complex queries into computational trees.
 Binary Tree Decomposition.
-For the one-to-many intersection/union structures in the tree, we separate each parent node into two child nodes.
-Note that the union branches merging step may create one-to-many structures that consist of both intersection and union edges, take Figure 1 for an example.
-This can be taken care of by first separating $v_{?}$ into an intersection structure ($v_{3}^{\prime}$ and $v_{5}^{\prime}$ in the example), and then separating the child node into an intersection structure ($v_{1}^{\prime}$ and $v_{2}^{\prime}$ for $v_{3}^{\prime}$, $v_{4}^{\prime},v_{3}$ for $v_{5}^{\prime}$), where $v^{\prime}$ denotes an intermediate entity retrieved by Neighborhood Retrieval Algorithm (Seen in Appendix A).
+We split each 1-to-n intersection/union node into n corresponding child nodes.
+Consider that merging union branches may result in a 1-to-n structure consisting of intersection and union edges, as shown in Figure 1.
+This can be properly addressed by separating $v_{?}$ into an intersection node structure ($v_{3}^{\prime}$ and $v_{5}^{\prime}$ in Figure 1) firstly, then decomposing the node that was decomposed in the previous layer into a deeper intersection node structure ($v_{1}^{\prime}$ and $v_{2}^{\prime}$ for $v_{3}^{\prime}$, $v_{4}^{\prime},v_{3}$ for $v_{5}^{\prime}$), where $v^{\prime}$ denotes an intermediate entity retrieved by Neighborhood Retrieval Algorithm (Seen in Appendix).
 Reverse Level Traversal.
@@
 Since the root node of the calculation tree is the answer entity, we perform a hierarchical traversal of all non-leaf nodes of the binary tree in reverse.
-As shown in Figure 1, the complex FOL query is decomposed into a sequence:
+As shown in Figure 1, the complex $EFO_{1}$ query is decomposed into a sequence:
 $[(v_{1},r,v_{1}^{\prime})$ $,(v_{2},r,v_{2}^{\prime}),(v_{3},r,v_{5}^{\prime}),(v_{4},r,v_{4}^{\prime}),$
@@
 Data Augmentation.
-Now we can turn any loopless FOL query into a series of separate subqueries.
+Now we can turn any loopless $EFO_{1}$ query into a series of separate subqueries.
 We use a defined template to integrate the decomposition process into the answers to the training corpus.
@@
 where $V_{\tau,\mathcal{D}ecomposed}$ indicates the answer corresponding to the logical query with the decomposition reasoning path.
-4.3 Fine-tuning Enhanced by Curriculum Learning
-As mentioned in previous sections, though decomposing into chain responses, complex queries still vary greatly in difficulty and complexity due to differences between query structure.
+Fine-tuning Enhanced by Curriculum Learning
+As mentioned in previous sections, though decomposing into chain responses, complex queries still vary greatly in difficulty and complexity due to differences between query structures.
 Naturally, we believe that these different types of samples should not be simply lumped together.
 Intuitively, we incorporate curriculum learning into our training.
-To be specific, In view of the particularity of complex reasoning data, when we decompose it into logical chains, naturally, we can use the number of decomposed sub-logical queries as a natural difficulty discriminator to select different types of queries, e.g., a 1p query would be defined as difficulty-1, while a 2p query, which can be decomposed into two projection queries and an intersection query, would be defined as difficulty-3.
-The detailed difficulty discriminating process will be shown in Appendix (Table S1).
+To be specific, given the particularity of complex reasoning data, when we decompose it into logical chains, naturally, we can use the number of decomposed sub-logical queries as a natural difficulty discriminator to select different types of queries, e.g., a 1p query would be defined as difficulty-1, while a 2p query, which can be decomposed into two projection queries and an intersection query, would be defined as difficulty-3.
+The detailed difficulty discriminating process will be shown in the Appendix.
 Finally, we divided samples into three parts: easy samples, medium samples and difficult samples according to the difficulty level.
@@
 On the contrary, we decided to first use 80% easy samples, 10% medium samples, and 10% difficult samples for the first stage of training and the subsequent two-stage training process is a Leto, and experimental results in the next few sections also proved that this is effective.
-Method $avg_{p}$ $avg_{ood}$ $avg_{n}$ 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni FB15K GQE 28.0 20.1 - 54.6 15.3 10.8 39.7 51.4 27.6 19.1 22.1 11.6 - - - - - Query2Box 38.0 29.3 - 68.0 21.0 14.2 55.1 66.5 39.4 26.1 35.1 16.7 - - - - - BetaE 41.6 34.3 11.8 65.1 25.7 24.7 55.8 66.5 43.9 28.1 40.1 25.2 14.3 14.7 11.5 6.5 12.4 CQD-CO 46.9 35.3 - 89.2 25.3 13.4 74.4 78.3 44.1 33.2 41.8 21.9 - - - - - CQD-Beam 58.2 49.8 - 89.2 54.3 28.6 74.4 78.3 58.2 67.7 42.4 30.9 - - - - - ConE 49.8 43.4 14.8 73.3 33.8 29.2 64.4 73.7 50.9 35.7 55.7 31.4 17.9 18.7 12.5 9.8 15.1 GNN-QE 72.8 68.9 38.6 88.5 69.3 58.7 79.7 83.5 69.9 70.4 74.1 61.0 44.7 41.7 42.0 30.1 34.3 QTO 74.0 71.8 49.2 89.5 67.4 58.8 80.3 83.6 75.2 74.0 76.7 61.3 61.1 61.2 47.6 48.9 27.5 LARK 56.1 43.1 18.4 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 82.6 71.9 56.9 93.5 73.5 59.6 92.3 82.3 76.8 75.9 74.6 60.4 81.2 61.6 52.0 43.5 41.7 FB15K-237 GQE 16.3 10.3 - 35.0 7.2 5.3 23.3 34.6 16.5 10.7 8.2 5.7 - - - - - Query2Box 20.1 15.7 - 40.6 9.4 6.8 29.5 42.3 21.2 12.6 11.3 7.6 - - - - - BetaE 20.9 14.3 5.5 39.0 10.9 10.0 28.8 42.5 22.4 12.6 12.4 9.7 5.1 7.9 7.4 3.5 3.4 CQD-CO 21.8 15.6 - 46.7 9.5 6.3 31.2 40.6 23.6 16.0 14.5 8.2 - - - - - CQD-Beam 22.3 15.7 - 46.7 11.6 8.0 31.2 40.6 21.2 18.7 14.6 8.4 - - - - - FuzzQE 24.0 17.4 7.8 42.8 12.9 10.3 33.3 46.9 26.9 17.8 14.6 10.3 8.5 11.6 7.8 5.2 5.8 ConE 23.4 16.2 5.9 41.8 12.8 11.0 32.6 47.3 25.5 14.0 14.5 10.8 5.4 8.6 7.8 4.0 3.6 GNN-QE 26.8 19.9 10.2 42.8 14.7 11.8 38.3 54.1 31.1 18.9 16.2 13.4 10.0 16.8 9.3 7.2 7.8 QTO 33.5 27.6 15.5 49.0 21.4 21.2 43.1 56.8 38.1 28.0 22.7 21.4 16.8 26.7 15.1 13.6 5.4 LARK 50.7 41.0 10.6 73.6 40.5 26.8 46.1 43.1 49.9 22.9 62.8 28.3 6.5 3.4 23.2 16.5 3.2 LACT 57.0 44.4 21.9 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 NELL995 GQE 18.6 12.5 - 32.8 11.9 9.6 27.5 35.2 18.4 14.4 8.5 8.8 - - - - - Query2Box 22.9 15.2 - 42.2 14.0 11.2 33.3 44.5 22.4 16.8 11.3 10.3 - - - - - BetaE 24.6 14.8 5.9 53.0 13.0 11.4 37.6 47.5 24.1 14.3 12.2 8.5 5.1 7.8 10.0 3.1 3.5 CQD-CO 28.8 20.7 - 60.4 17.8 12.7 39.3 46.6 30.1 22.0 17.3 13.2 - - - - - CQD-Beam 28.6 19.8 - 60.4 20.6 11.6 39.3 46.6 25.4 23.9 17.5 12.2 - - - - - FuzzQE 27.0 18.4 7.8 47.4 17.2 14.6 39.5 49.2 26.2 20.6 15.3 12.6 7.8 9.8 11.1 4.9 5.5 ConE 27.2 17.6 6.4 53.1 16.1 13.9 40.0 50.8 26.3 17.5 15.3 11.3 5.7 8.1 10.8 3.5 3.9 GNN-QE 28.9 19.6 9.7 53.3 18.9 14.9 42.4 52.5 30.8 18.9 15.9 12.6 9.9 14.6 11.4 6.3 6.3 QTO 32.9 24.0 12.9 60.7 24.1 21.6 42.5 50.6 31.3 26.5 20.4 17.9 13.8 17.9 16.9 9.9 5.9 LARK 52.9 26.9 12.4 87.8 45.7 33.5 51.3 48.7 23.1 22.2 20.6 41.1 9.9 5.9 24.5 13.3 7.3 LACT 60.1 32.0 17.2 91.4 53.6 40.6 62.2 54.9 31.4 34.8 27.0 34.0 16.0 21.2 21.0 16.3 11.6
-Table 1: Test MRR results (%) on complex query answering across all query types.
-$avg_{p}$ is the average on EPFO queries; $avg_{ood}$ is the average on out-of-distribution (OOD) queries, which consist pi/ip/2u/ip queries; $avg_{n}$ is the average on queries with negation.
-4.4 Reasoning Module
-We use the final LACT LLM as the answer generator, as shown in Figure 1(c).
-We retrieve relevant information and textualize the FOL query, and finally we populate it into the template in 4.4 to generate responses.
-We use the LLM to do a simple text generation task to get the answer.
-After fine-tune, LACT LLM can follow the output mode in training stage in Figure 1, so we can extract final answers through simple regular expressions with the template in 4.4.
-Prompt 1: Query Prompt Template of LACT. [Uncaptioned image]
-Prompt 2: Answer Template of LACT. [Uncaptioned image]
-5 Experiments
-5.1 Training Settings
-Training datasets We opt for the the most popular datasets: FB15K, FB15K-237, NELL995.
-Detailed information about dataset is listed in  D.1.
+Method $avg_{p}$ $avg_{ood}$ $avg_{n}$ 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Dataset FB15K GQE 28.0 20.1 - 54.6 15.3 10.8 39.7 51.4 27.6 19.1 22.1 11.6 - - - - - Query2Box 38.0 29.3 - 68.0 21.0 14.2 55.1 66.5 39.4 26.1 35.1 16.7 - - - - - BetaE 41.6 34.3 11.8 65.1 25.7 24.7 55.8 66.5 43.9 28.1 40.1 25.2 14.3 14.7 11.5 6.5 12.4 CQD-CO 46.9 35.3 - 89.2 25.3 13.4 74.4 78.3 44.1 33.2 41.8 21.9 - - - - - CQD-Beam 58.2 49.8 - 89.2 54.3 28.6 74.4 78.3 58.2 67.7 42.4 30.9 - - - - - ConE 49.8 43.4 14.8 73.3 33.8 29.2 64.4 73.7 50.9 35.7 55.7 31.4 17.9 18.7 12.5 9.8 15.1 GNN-QE 72.8 68.9 38.6 88.5 69.3 58.7 79.7 83.5 69.9 70.4 74.1 61.0 44.7 41.7 42.0 30.1 34.3 QTO 74.0 71.8 49.2 89.5 67.4 58.8 80.3 83.6 75.2 74.0 76.7 61.3 61.1 61.2 47.6 48.9 27.5 LARK 56.1 43.1 18.4 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 82.6 71.9 56.9 93.5 73.5 59.6 92.3 82.3 76.8 75.9 74.6 60.4 81.2 61.6 52.0 43.5 41.7 Dataset FB15K-237 GQE 16.3 10.3 - 35.0 7.2 5.3 23.3 34.6 16.5 10.7 8.2 5.7 - - - - - Query2Box 20.1 15.7 - 40.6 9.4 6.8 29.5 42.3 21.2 12.6 11.3 7.6 - - - - - BetaE 20.9 14.3 5.5 39.0 10.9 10.0 28.8 42.5 22.4 12.6 12.4 9.7 5.1 7.9 7.4 3.5 3.4 CQD-CO 21.8 15.6 - 46.7 9.5 6.3 31.2 40.6 23.6 16.0 14.5 8.2 - - - - - CQD-Beam 22.3 15.7 - 46.7 11.6 8.0 31.2 40.6 21.2 18.7 14.6 8.4 - - - - - FuzzQE 24.0 17.4 7.8 42.8 12.9 10.3 33.3 46.9 26.9 17.8 14.6 10.3 8.5 11.6 7.8 5.2 5.8 ConE 23.4 16.2 5.9 41.8 12.8 11.0 32.6 47.3 25.5 14.0 14.5 10.8 5.4 8.6 7.8 4.0 3.6 GNN-QE 26.8 19.9 10.2 42.8 14.7 11.8 38.3 54.1 31.1 18.9 16.2 13.4 10.0 16.8 9.3 7.2 7.8 QTO 33.5 27.6 15.5 49.0 21.4 21.2 43.1 56.8 38.1 28.0 22.7 21.4 16.8 26.7 15.1 13.6 5.4 LARK 50.7 41.0 10.6 73.6 40.5 26.8 46.1 43.1 49.9 22.9 62.8 28.3 6.5 3.4 23.2 16.5 3.2 LACT 57.0 44.4 21.9 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 Dataset NELL995 GQE 18.6 12.5 - 32.8 11.9 9.6 27.5 35.2 18.4 14.4 8.5 8.8 - - - - - Query2Box 22.9 15.2 - 42.2 14.0 11.2 33.3 44.5 22.4 16.8 11.3 10.3 - - - - - BetaE 24.6 14.8 5.9 53.0 13.0 11.4 37.6 47.5 24.1 14.3 12.2 8.5 5.1 7.8 10.0 3.1 3.5 CQD-CO 28.8 20.7 - 60.4 17.8 12.7 39.3 46.6 30.1 22.0 17.3 13.2 - - - - - CQD-Beam 28.6 19.8 - 60.4 20.6 11.6 39.3 46.6 25.4 23.9 17.5 12.2 - - - - - FuzzQE 27.0 18.4 7.8 47.4 17.2 14.6 39.5 49.2 26.2 20.6 15.3 12.6 7.8 9.8 11.1 4.9 5.5 ConE 27.2 17.6 6.4 53.1 16.1 13.9 40.0 50.8 26.3 17.5 15.3 11.3 5.7 8.1 10.8 3.5 3.9 GNN-QE 28.9 19.6 9.7 53.3 18.9 14.9 42.4 52.5 30.8 18.9 15.9 12.6 9.9 14.6 11.4 6.3 6.3 QTO 32.9 24.0 12.9 60.7 24.1 21.6 42.5 50.6 31.3 26.5 20.4 17.9 13.8 17.9 16.9 9.9 5.9 LARK 52.9 26.9 12.4 87.8 45.7 33.5 51.3 48.7 23.1 22.2 20.6 41.1 9.9 5.9 24.5 13.3 7.3 LACT 60.1 32.0 17.2 91.4 53.6 40.6 62.2 54.9 31.4 34.8 27.0 34.0 16.0 21.2 21.0 16.3 11.6
+Table 1: MRR scores (%) on complex reasoning for 14 types of queries.
+$avg_{p}$ represents the mean score of nine ordinary queries including 1p/2p/3p/2i/3i and pi/ip/2u/ip; $avg_{ood}$ is the mean score of out of distribution (OOD) queries, which consist pi/ip/2u/ip queries; $avg_{n}$ is the mean score of queries including negation operation.
+Reasoning Module
+We retrieve relevant information and textualize the $EFO_{1}$ query, and finally, we populate it into the template in the Appendix to generate queries.
+As shown in Appendix, we use the LLM to do a simple text generation task to get the answer.
+After fine-tuning, LACT can follow the output mode in the training stage in Figure 1, so we can extract final answers through simple regular expressions with the template in the Appendix.
+The specific prompt can be found in the Appendix.
+Experiments
+Training Settings
+We opt for the most popular datasets: FB15K, FB15K-237, NELL995.
 We used the training set of the above dataset as the original training data.
-Training Details We use open-source model LLaMA-2-base, including two different parameter sizes: 7B and 13B, as the base model for fine-tuning.
-All LLaMA-2-7B and LLaMA-2-13B models are trained by fully fine-tuning.
-For the fully fine-tuning setting, we use the AdamW optimizer to train the model with 1 epoch and the batch size is 128. We use 8 NVIDIA A100 GPUS for training with the learning rate of 3e-6.
-Training Cost We trained for a total of approximately 10 hours using 8*A100 during three training stages.
-5.2 Experimental Settings
-Baseline Methods For comparing with KGE, we chose the following representative methods as baselines: GQE (?), Query2Box(Q2B) (?), BetaE (?), CQD (?), ConE (?), GNN-QE (?), QTO (?).
-We also compared our method to an LLM-based method, LARK  (?).
-Evaluation Protocol Following previous works (?), we use mean reciprocal rank (MRR) as standard evaluation protocols for evaluating complex reasoning over knowledge graph.
-In the filtered setting, all easy and hard answers are filtered out during ranking.
-The detail could be found in Appendix  D.3.
-5.3 Main Results
+Detailed information about the dataset and training set is listed in the Appendix.
+Experimental Settings
+Baseline Methods For comparing with KGE, we chose the following representative methods as baselines: GQE (Hamilton et al.
+2018), Query2Box(Q2B) (Choudhary and Reddy 2023), BetaE (Ren and Leskovec 2020), CQD (Arakelyan et al.
+2020), ConE (Zhang et al.
+2021), GNN-QE (Zhu et al.
+2022), QTO (Bai et al.
+2023).
+Moreover, we also compared our LACT with the LLM-based method LARK  (Choudhary and Reddy 2023).
+Figure 2: Performance of scaling LACT on FB15K-237 with different a) model and b) data scales.
+Figure 3: Results of Ablation Studies.
+(a) Comparison PPL and train loss results of whether to use BTD based on FB15k.
+(b) Comparison PPL results of whether use CL based on FB15K high-difficulty queries.
+Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2 67.2 42.3 38.3 61.6 44.8 34.1 36.9 44.2 28.4 44.7 38.5 36.9 32.1 30.0    + IT 94.6 68.8 60.2 84.5 66.7 56.0 60.2 69.5 42.3 66.5 54.4 41.0 38.8 36.9     +BTD 91.5 72.3 65.6 89.7 75.2 60.1 65.4 72.5 49.9 74.3 66.4 48.9 46.5 42.5
+Table 2: Accuracy results (%) evaluated on FB15k of whether to use BTD for hard complex query on all query types.
+Evaluation Protocol Following previous works (Ren, Hu, and Leskovec 2019), we choose the Mean Reciprocal Rank (MRR) with standard evaluation protocol for evaluating complex reasoning on KG.
+The detailed setting can be found in the Appendix.
+Main Results
 The main experiment results of three datasets are shown in Table 1.
-The baselines were trained on 1p/2p/3p/2i/3i queries, hence other than these 4 types of EPFO queries serve as OOD queries, and we report the average result on these queries in $avg_{ood}$.
-We observe that LACT significantly outperforms baseline methods across all datasets.
-Notably, LACT yields an averag gain of 7.3%, 2.9%, and 6.3% on $avg_{p}$, $avg_{odd}$, and $avg_{n}$, compared to the previous SOTA method, especially more challenging datasets like FB15K-237 and NELL995.
-This suggests that our method has better reasoning capability and captures a broad range of relations to effectively utilize this capability for enhancing the performance of complex queries.
-Figure 2: Results of Ablation Studies.(a) Comparison PPL and train loss results of whether to use BTD based on FB15k.(b) Comparison PPL results of whether use CL based on FB15K high-difficulty queries.
-Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2 67.2 42.3 38.3 61.6 44.8 34.1 36.9 44.2 28.4 44.7 38.5 36.9 32.1 30.0    + IT 94.6 68.8 60.2 84.5 66.7 56.0 60.2 69.5 42.3 66.5 54.4 41.0 38.8 36.9     +BTD 91.5 72.3 65.6 89.7 75.2 60.1 65.4 72.5 49.9 74.3 66.4 48.9 46.5 42.5
-Table 2: Accuracy results (%) of whether to use BTD on hard complex query answering across all query types, evaluated on FB15k.
+The baseline training datasets contain five different queries including 1p/2p/3p/2i/3i, so another four queries are considered as OOD types, so the mean value of OOD types of queries is recorded as $avg_{ood}$.
+In our experiments, LACT consistently outperforms baseline methods across all datasets.
+Notably, LACT yields an average gain of 7.3%, 2.9%, and 6.3% on $avg_{p}$, $avg_{odd}$, and $avg_{n}$, compared to the previous SOTA method, especially more challenging datasets like FB15K-237 and NELL995.
+Our method exhibits superior reasoning capability and effectively captures a wide range of internal relations, leading to enhanced performance on complex queries.
 BTD CL 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2 w/IT 94.6 68.8 60.2 84.5 66.7 56.0 60.2 69.5 42.3 66.5 54.4 41.0 38.8 36.9 94.2 72.3 65.6 89.7 75.2 60.1 65.4 72.5 49.9 74.3 66.4 48.9 46.5 42.5 94.6 70.9 61.3 84.4 72.7 58.2 63.2 69.5 47.3 68.5 64.0 42.3 40.9 37.1 94.8 78.7 69.2 94.8 88.1 79.3 80.5 80.7 67.1 90.6 70.4 59.3 53.6 46.7
-Table 3: Accuracy results (%) of whether to use CL on hard complex query answering across all query types, evaluated on FB15k.
-Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2-7B 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 Mistral-7B 82.6 59.5 34.2 59.6 58.5 59.1 39.1 59.2 33.7 20.4 39.2 31.4 22.4 13.5 Qwen1.5-7B 81.3 57.2 36.8 62.9 58.1 63.9 38.6 60.4 31.8 20.7 41.9 28.7 20.8 11.6
-Table 4: MRR results (%) based on differnet base models across all query types, evaluated on FB15k-237.
-Figure 3: The results of the main experiment. We evaluate the performance of three current state-of-the-art methods on three datasets.
-5.4 Ablation Studies
-To verify the effectiveness of the LACT design, we conduct a two-part ablation study.
-The first part is designed to verify the effectiveness of logical chain decomposition and the second part is designed to verify the effectiveness of curriculum learning.
+Table 3: Accuracy results (%) evaluated on FB15k of whether to use CL for hard complex query on all query types.
+Ablation Studies
+To validate the efficacy of the LACT module, we conduct a two-part ablation study.
+Firstly, we will investigate the impact of logical chain decomposition, while the second part assesses the effectiveness of curriculum learning.”
 Effect of Binary Tree Decomposition (BTD).
 As shown in Figure 3, logical chain decomposition can stimulate LLM’s ability of logical decomposition, thereby greatly improving the performance of difficult queries.
-From a training perspective, as shown in Figure 2, although perplexity (PPL) and training loss of decomposed queries before training was slightly higher than that of ordinary queries, we found that as training progresses, the loss and PPL of decomposed queries will quickly decrease to levels much lower than ordinary queries, proving that chain decomposition is effective to reduce the difficulty of learning complex queries.
+From a training perspective, as shown in Figure 3, although perplexity (PPL) and training loss of decomposed queries before training were slightly higher than that of ordinary queries, we found that as training progresses, the loss and PPL of decomposed queries will quickly decrease to levels much lower than ordinary queries, proving that chain decomposition is effective to reduce the difficulty of learning complex queries.
 Effect of Curriculum Learning. Curriculum learning, as illustrated in Table 3, greatly alleviates the gap between difficult training samples and the understanding ability of LLMs.
@@
 It is difficult for LLMs to understand the difficulty difference between undecomposed samples, so curriculum learning is also difficult to take effect.
-Figure 4: Performance of scaling LACT on FB15K-237 with different a) model and b) data scales.
-5.5 Transferability Study
+Transferability Study
 Considering the diversity of complex reasoning tasks, we can divide transferability into two levels, task-level and dataset-level transferability.
 Task-level transferability.
-The results in Table 1 show that our method achieves a relative gain of 9.9% on the OOD task, which demonstrates the strong generalization of our fine-tuning framework.
-Even in the OOD queries, as shown in Table 5, more than 95% of test samples can still follow logical chain reasoning.
-These phenomena indicate strong generalization ability of LACT.
-Metric pi ip 2u up Prodecomposed 98.7 100.0 97.8 100.0 Protrue,decomposed 98.6 99.9 97.8 99.6
-Table 5: In OOD queries, the proportion of queries that can be decomposed and the proportion of queries that can be decomposed correctly on fb15k.
-Figure 5: Ablation experimental results of Accuracy (%) trained on FB15k and tested on FB15K-237, compared to models trained on all mixed training data.
+The results in Table 1 show that our LACT achieves a relative gain of 9.9% on the OOD task, which demonstrates the strong generalization of our fine-tuning framework.
+Even in the OOD queries, as shown in the Appendix, more than 95% of test samples can still follow logical chain reasoning.
+These phenomena indicate the strong generalization ability of LACT.
 Dataset-level transferability.
 In fact, almost all KGE methods, even if some of the optimization methods claim not to require training, require a KGE base model adapted to a specific dataset, which leads to the inherent defect of extremely poor Transferability of the KGE method.
-However, as previous research has shown, fine-tuning of LLMs is mainly to stimulate the knowledge and capabilities of potentially relevant downstream tasks contained in LLM pre-training.
+However, as previous research has shown, fine-tuning LLMs is mainly to stimulate the knowledge and capabilities of potentially relevant downstream tasks contained in LLM pre-training.
 This has also become the theoretical basis for the transferability of fine-tuning methods for LLMs.
-The results in Figure 5 show that the reasoning ability stimulated by one dataset can still be demonstrated in another dataset, which reflects well in the query performance which only dropped less than 5%.
+The results in the Appendix show that the reasoning ability stimulated by one dataset can still be demonstrated in another dataset, which reflects well in the query performance which only dropped less than 5%.
 Model-level transferability.
 We tried analytical experiments with different base models to determine whether our LACT was universal.
-Obviously, all types of queries have been improved to a certain extent due to the progress of the base model.
+Obviously, The results in the Appendix show that all types of queries have been improved to a certain extent due to the progress of the base model.
 Experimental results show that our LACT is suitable for different pedestal models and has strong generalization.
-5.6 Scalability Study
-For verifying the scalability of LACT, we scale LACT to different model sizes and data volumes.
+Scalability Study
+To verify the scalability of LACT, we scale LACT to different model sizes and data volumes.
 Performance on different model size.
@@
 Figure 4: (left) MRR performance of LACT and previous SOTA methods at different difficulties based on FB15K.
-(right) The correlation between related information completeness and accuracy evaluated on FB15K-237, we selected 3p query and pi query with the same inference path length as the task types.
+(right) The correlation between related information completeness and accuracy evaluated on FB15K-237, we selected a 3p query and pi query with the same inference path length as the task types.
 We assume that the completeness of all simple queries is 1.
-6 Discussion
-Figure 7: Inference results of ChatGPT and LACT on 2i query case, respectively.
-6.1 When and Where Does LACT Work?
+Discussion
+When and Where Does LACT Work?
 The performance of LACT would be related to the following two aspects: I. The completeness of relevant information extracted from KG. II. Sophistication of complex reasoning.
@@
 We take the form of a posteriori that set the completeness of relevant triplets to the proportion of triplets in the inference path of complex reasoning in the provided context, (For example, if the inference path of a 2i query is (Turing Award, winner, Yoshua Bengio), (Canada, citizen, Yoshua Bengio), and we can only retrieve (Turing Award, winner, Yoshua Bengio) in incomplete KG, then our completeness of relevant triplets to 1/2.) and set the completeness of simple queries that can be directly inferred to 1, to obtain the relation between Accuracy and correlation information completeness.
-As seen in Figure  6, LACT obtains a significant gain when the completeness of relevant information increased, though, with zero relevant information, it remains a certain amount of complex reasoning ability.
-LACT performs consistently better on higher difficulties As mentioned before, we simply divide the difficulty of the query by the number of hops in the query.
-The results in Figure  6 show that our model yields more gain in tasks of higher-level difficulty and complexity, which benefits from our unique and sophisticated fine-tuning framework.
-6.2 Case Study
-To have a close look, we perform the case studies by analyzing the results of LACT and ChatGPT (GPT-3.5-turbo-0613).
-As shown in Figure 7, ChatGPT cannot make good use of incomplete knowledge graphs for reasoning in some cases.
-Conversely, LACT performs reasoning through a complete logical chain, making maximum use of the relevant information provided and deducing the correct answer, which greatly improves the reasoning ability.
-7 Conclusion
+As seen in Figure  4, LACT obtains a significant gain when the completeness of relevant information increases, though, with zero relevant information, it remains a certain amount of complex reasoning ability.
+LACT performs consistently better on higher difficulties.
+As mentioned before, we simply divide the difficulty of the query by the number of hops in the query.
+The results in Figure  4 show that our model yields more gain in tasks of higher-level difficulty and complexity, which benefits from our unique and sophisticated fine-tuning framework.
+Conclusion
 In this paper, we present a simple and effective fine-tuning framework LACT to boost the complex logical reasoning ability over KGs.
-LACT is a two-stage method that consists of both Binary Tree Decomposition and Curriculum Learning and can be applied to various size LLMs with different data sizes.
-We empirically demonstrate the effectiveness and universality of the LACT on a series of widely-used knowledge graph datasets.
-Further analyses reveal the underlying mechanism of our method, and investigate When and Why Does LACT Work.
+LACT is a two-part method that consists of both Binary Tree Decomposition and Curriculum Learning and can be applied to various size LLMs with different data sizes.
+We empirically demonstrate the effectiveness and universality of the LACT on a series of widely used knowledge graph datasets.
+Further analyses delve into the underlying mechanism of our LACT, and investigate When and Why Does LACT Work.
 We hope that our work can inspire more research on combining knowledge graphs and LLMs.
+Acknowledgments
+This research is supported by NSFC (62225113, 62206202), National Key RD Program of China (2023YFC2705700), The Innovative Research Group Project of Hubei Province under Grants (2024AFA017), the National Research Foundation, Singapore, and the CyberSG R&D Programme Office (“CRPO”), under the National Cybersecurity R&D Programme (“NCRP”), RIE2025 NCRP Funding Initiative (Award CRPO-GC1-NTU-002).
 Appendix
 Appendix A Neighborhood Retrieval Algorithm
-A.1 Retrieval Algorithm
+Retrieval Algorithm
 To strike a balance between the completeness of relevant information and the token number limit of LLMs, we search for as many relevant triplets as possible along the possible paths.
 Particularly, for the 1p query, we simply find all the triplets containing the entity or the relation.
-For another query, as shown in Figure S1 for each leaf node in DAG, we do depth traversal on the graph.
+For another query, as shown in Figure 6 for each leaf node in DAG, we do depth traversal on the graph.
 For each step in the traversal process, if this step is a projection, we search for all the possible triplets.
@@
 We continue this traversal until the obtained entity is empty or reaches the root node. All triplets during the traversal are related to triplets.
-A.2 Over-limit Solutions and Token distribution
+Over-limit Solutions and Token distribution
 In all experimental setups, we used the Max Seq Length of 4096, and we simply discarded all out-of-bounds samples and recorded them as 0 at the time of evaluation.
 In fact, after using our information retrieval algorithm, most of our samples were controlled below 4096.
-As shown in Figure S2, compared with the algorithm of searching all possible related triples, our retrieval algorithm greatly reduces the consumption of tokens.
-Figure S1: A case on Neighborhood Retrieval Algorithm
-Figure S2: Probability distribution of the number of tokens in FB15K and all-context Retrieval Algorithm.
+As shown in Figure 7, compared with the algorithm of searching all possible related triples, our retrieval algorithm greatly reduces the consumption of tokens.
+Figure 5: Query structure diagram of 14 types of queries
+Figure 6: A case on Neighborhood Retrieval Algorithm
+Figure 7: Token numbers’ probability distribution in the dataset FB15K-237 and all-context Retrieval Algorithm.
 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Number of subqueries 1 2 3 3 5 4 4 3 4 3 5 4 4 4 Difficulty 1 1 2 2 3 3 3 2 3 2 3 3 3 3
-Table S1: Difficulty of different query types where 1 means easy, 2 means medium, 3 means hard.
-Figure S3: Query structures, illustrated in their query computation graph representations
-Appendix B Conversion Between FOL Expression and Query Computation Tree
-The conversion from an FOL expression (disjunctive normal form) to its query computation tree involves two steps: dependency graph generation, and union branches duplication.
+Table 4: Difficulty of different query types where 1 means easy, 2 means medium, 3 means hard.
+Appendix B Universal Procedure for Converting from $EFO_{1}$ to Computational Trees
+To transform a $EFO_{1}$ formula into its corresponding computation tree, a two-phase process is employed: dependency graph generation, and union branches duplication.
 Dependency Graph Generation.
-Upon encountering a First-Order Logic (FOL) expression, our primary procedure entails the allocation of distinct nodes to individual variables, while assigning a unique node to the constant entity within each one-hop atom.
-It is important to acknowledge that multiple nodes may represent the same constant entity, given its occurrence in various one-hop atoms.
+Upon encountering a $EFO_{1}$ expression, our primary procedure entails the allocation of distinct nodes to individual variables, while a separate node is assigned to each specific entity for each single-hop query.
+It is important to acknowledge that multiple nodes may represent one same intermediate entity, given its occurrence among various single-hop atom.
 Subsequently, undirected edges are employed to establish connections between nodes in accordance with the defined one-hop atoms.
-Specifically, if $e^{i}_{j}=r(v^{\prime},v)$ (or $r(c,v)$), then we connect the nodes of $v^{\prime}$ (or $c$) and $v$ by an edge $r_{i}$.
-Similarly, if $e^{i}_{j}=\lnot r(v^{\prime},v)$ (or $\lnot r(c,v)$), then we connect the nodes of $v^{\prime}$ (or $c$) and $v$ by an edge $\lnot r_{i}$.
-The variable $i$ serves as a distinguishing label for edges emanating from distinct conjunctions.
-The formulated undirected dependency multigraph must conform to a tree structure, signifying a connected and acyclic graph.
-Choosing the node $v_{?}$ as the root, we establish edge directions ensuring that they uniformly point from child nodes to their respective parent nodes, with due consideration to handling inverse relations.
-Notably, constant entities inherently function as leaf nodes within the tree, given that each entity node exclusively connects to a single variable node.
-Union Branches Duplication Then we handle the duplication branches in the query computation tree.
-On the path $\tau$ from root to every leaf node, if exists, we find the first node $v_{i}$ such that the edges between $v_{i}$ and its child node $v_{j}$ are all of the same relations, but in different conjunctions: $r_{t_{1}},r_{t_{2}},\dots,r_{t_{p}}$.
-We merge these edges into a single edge $r_{t_{1},t_{2},\dots t_{p}}$, since they all correspond to the same one-hop atom but in different conjunctions, they can be merged by the distributive law:
+In the case where $e^{i}_{j}=r(h,t)$, we link $h$ and $t$ by establishing the edge recorded as $h_{k}$.
+Similarly, if $e^{i}_{j}=\lnot r(h,t)$, we link $h$ and $t$ by the edge recorded as $\lnot r_{k}$.
+The variable $k$ serves as a distinguishing label for edges emanating from distinct atom struture.
+The formulated multi-graph must conform be in a tree-shape, signifying a DAG.
+Choosing tail node $v_{?}$ to be the root, We define the orientation of edges to consistently direct from child nodes towards their parent nodes, while carefully accommodating inverse relationships.
+It is important to note that constant entities nherently serve as terminal nodes in the computation tree, as they are each uniquely associated with a solitary variable node.
+Union Branches Duplication Subsequently we address the duplication branches within the computation tree.
+For each path $\tau$ from every leaf node to root, we search for the first node $v_{i}$ which contains the same relations between $v_{i}$ and its possible child nodes $v_{i}^{{}^{\prime}}$, but in different link structure: $r_{i_{1}},r_{i_{2}},\dots,r_{i_{n}}$.
+These edges were amalgamate into a converged edge $r_{i_{1},i_{2},\dots i_{n}}$ to dissolve the complex multiple edges structure and then they can be amalgamate abide by the following distributive law:
 $(A\land B)\lor(A\land C)\Leftrightarrow A\land(B\lor C)$ (7)
-We assert that there is a subpath from $v_{i}$ to some $v_{k}$ within the path $\tau$ that only consists of edges $r_{t_{1},t_{2},\dots t_{p}}$, and $v_{k}$ is connected to different child nodes by relations from conjunctions $t_{1},t_{2},\dots t_{p}$.
-These edges are annotated as unions, while the remaining one-to-many structures are designated as intersections.
-Consequently, the multigraph transforms into a simple graph, devoid of multiple edges.
-Appendix C Difficulty
-We divide the difficulty by the number of decomposed subqueries. Query types and their corresponding difficulties are shown in Table S1.
-Appendix D Experiment Details
-D.1 Dataset Details
-$\bullet$ FB15K is based on Freebase, a large collaborative knowledge graph project that was created by Google.
-FB15k contains about 15,000 entities, 1,345 relations, and 592,213 triplets (statements that assert a fact about an entity).
-$\bullet$ FB15K-237 is a subset of FB15k, containing 14,541 entities, 237 relations, and 310,116 triplets.
-The relations in FB15k-237 are a subset of the relations in FB15k, and were created to address some of the limitations of FB15k, such as the presence of many irrelevant or ambiguous relations, and to provide a more challenging benchmark for knowledge graph completion models.
-$\bullet$ NELL995 was created using the Never-Ending Language Learning (NELL) system, which is a machine learning system that automatically extracts knowledge from the web by reading text and inferring new facts.
-NELL995 contains 9,959 entities, 200 relations, and 114,934 triplets.
-The relations in NELL995 cover a wide range of domains, including geography, sports, and politics.
-D.2 Query Structure
-For a fair comparison, we use the 14 types of complex queries generated by the same rules in (?). The query structure of each type is shown in Figure S3.
-For each complex query, answers are divided into easy answers and hard answers, based on whether answer can be derived by existing edges in the graph directly.
-Specifically, in the valid/test set, the easy answers are the entities that can be inferred by edges in the training/valid graph, while hard answers are those that need to be inferred by predicting missing edges in the valid/test graph.
-Referring to previous work, we calculate standard evaluation metrics including mean reciprocal rank (MRR), in the filtered setting where all easy and hard answers are filtered out during ranking.
-D.3 Evaluation Protocol Detail
-For each complex query, its answers are divided into easy answers and hard answers, based on whether the answer can be derived by existing edges in the graph directly.
-Specifically, in the valid/test set, the easy answers are the entities that can be inferred by edges in the training/valid graph, while hard answers are those that need to be inferred by predicting missing edges in the valid/test graph.
-Referring to previous work (?), we calculate standard evaluation metrics including mean reciprocal rank (MRR), in the filtered setting where all easy and hard answers are filtered out during ranking.
+Appendix C Prompt Template
+The query and output prompt templates can be found in C and C.
+Prompt 1: Query Prompt Template of LACT. [Uncaptioned image]
+Prompt 2: Answer Template of LACT. [Uncaptioned image]
+Appendix D Difficulty
+We divide the difficulty by the number of decomposed subqueries. Query types and their corresponding difficulties are shown in Table 4.
+Appendix E Experiment Details
+Dataset Details
+$\bullet$ FB15K comes from a subset of Freebase, a large knowledge database initiated by Google. FB15k has a total of 592,213 triplets consisting of 15,000 entities and 1,345 relationships.
+$\bullet$ FB15K-237 is the subset of FB15k which contains 310,116 triplets consisting of 14,541 entities and 237 relations.
+FB15K-237 corrects the inverse relationship leakage problem in the test set to a certain extent by selecting a specific relationship subset in FB15K, thereby more realistically reflecting the effect of the model on complex reasoning on the knowledge graph.
+$\bullet$ NELL995 is automatically extracted and generated by Carnegie Mellon University’s Never-Ending Language Learning (NELL) project.
+It was considered as a comprehensive knowledge graph data set that includes 995 relationships in multiple fields.
+Training Details
+We use open-source model LLaMA-2-base, including two different parameter sizes: 7B and 13B, as the base model for fine-tuning. All LLaMA-2-7B and LLaMA-2-13B models are trained by fully fine-tuning.
+For the fully fine-tuning setting, we use the AdamW optimizer to train the model with 1 epoch and the batch size is 128. We use 8 NVIDIA A100 GPUS for training with the learning rate of 3e-6.
+Training Cost
+We trained for a total of approximately 10 hours using 8*A100 during three training stages.
+Query Structure
+Taking into account the fairness of the evaluation, we use the same 14 types of queries as in the previous work (Bai et al. 2023). The query structure of each type is shown in Figure 5.
+Figure 8: The results of the main experiment. We evaluate the performance of three current state-of-the-art methods on three datasets.
+Method 1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni Llama2-7B 76.5 54.3 30.3 56.0 54.5 54.6 36.9 56.5 29.7 17.6 33.1 27.1 19.8 11.2 Mistral-7B 82.6 59.5 34.2 59.6 58.5 59.1 39.1 59.2 33.7 20.4 39.2 31.4 22.4 13.5 Qwen1.5-7B 81.3 57.2 36.8 62.9 58.1 63.9 38.6 60.4 31.8 20.7 41.9 28.7 20.8 11.6
+Table 5: MRR results (%) based on differnet base models evaluated on FB15k-237.
+Evaluation Protocol Detail
+The answers to queries in complex logical reasoning can be divided into simple and hard types, which are distinguished by whether answers of the query can be easily retrieved directly in the knowledge graph.
+Specifically, in the valid set, we consider answers that can be directly retrieved from the graph in the training set as simple answers, and those that cannot be directly retrieved as hard answers.
+And in the test set, we consider answers that can be directly retrieved from the valid graph as simple answers, and those that cannot be directly retrieved as hard answers.
+Referring to previous work (Ren and Leskovec 2020), we choose the Mean Reciprocal Rank (MRR) as the evaluation index on the filtered setting all answers (including easy and hard) would be filtered out when ranking.
+Figure 9: Inference results of ChatGPT and LACT on 2i query case, respectively.
+Figure 10: Ablation experimental results of Accuracy (%) trained on FB15k and tested on FB15K-237, compared to models trained on all mixed training data.
+Appendix F More Results
+Overall comparison
+We provide a multi-dimensional comparison with previous state-of-the-art methods in Figure  8
+OOD results
+1p 2p 3p 2i 3i pi ip 2u up 2in 3in inp pin pni LARK 72.8 50.7 36.2 66.9 60.4 56.1 23.5 52.4 40.6 16.2 5.7 33.7 26.1 10.0 LACT 97.0 87.2 80.8 95.8 85.3 71.7 77.4 85.7 62.6 88.5 81.4 77.8 68.9 65.7
+Table 6: Accuracy results (%) evaluated on FB15k on complete KG for complex query on all query types.
+Metric pi ip 2u up nin nipn Prodecomposed 98.7 100.0 97.8 100.0 98.4 97.2 Protrue,decomposed 98.6 99.9 97.8 99.6 96.1 96.9
+Table 7: In OOD queries, the proportion of queries that can be decomposed and the proportion of queries that can be decomposed correctly on FB15k.
+As shown in Table 7including pi/ip/2u/up more than 95% of OOD test samples can still follow logical chain reasoning.
+These phenomena indicate strong generalization ability of LACT.
+In fact, many query types including nin/nipn that have not been tested before can also be correctly decomposed and queried by LACT.
+Dataset Ablation Results
+The results in Figure 10 show that the reasoning ability stimulated by one dataset can still be demonstrated in another dataset, which reflects well in the query performance which only dropped less than 5%.
+Model-level transferability
+The results in Table  5 show that LACT is effective on different base models and benefits from the improvement of the base model.
+Case Study
+To have a close and interen look, we perform the case studies by analyzing the results of LACT and ChatGPT (GPT-3.5-turbo-0613).
+As shown in Figure 9, ChatGPT cannot make good use of incomplete knowledge graphs for reasoning in some cases.
+Conversely, LACT performs reasoning through a complete logical chain, making maximum use of the relevant information provided and deducing the correct answer, which greatly improves the reasoning ability.